### v1

Abstract
We analyze various formulations of the $\infty$-Laplacian eigenvalue problem on graphs, comparing their properties and highlighting their respective advantages and limitations. First, we investigate the graph $\infty$-eigenpairs arising as limits of $p$-Laplacian eigenpairs, extending key results from the continuous setting to the discrete domain. We prove that every limit of $p$-Laplacian eigenpair, for $p$ going to $\infty$, satisfies a limit eigenvalue equation and establish that the corresponding eigenvalue can be bounded from below by the packing radius of the graph, indexed by the number of nodal domains induced by the eigenfunction. Additionally, we show that the limits, for $p$ going to $\infty$, of the variational $p$-Laplacian eigenvalues are bounded both from above and from below by the packing radii, achieving equality for the smallest two variational eigenvalues and corresponding packing radii of the graph. In the second part of the paper, we introduce generalized $\infty$-Laplacian eigenpairs as generalized critical points and values of the $\infty$-Rayleigh quotient. We prove that the generalized variational $\infty$-eigenvalues satisfy the same upper bounds in terms of packing radii as the limit of the variational eigenvalues, again with equality holding between the smallest two $\infty$-variational eigenvalues and the first and second packing radii of the graph. Moreover, we establish that any solution to the limit eigenvalue equation is also a generalized eigenpair, while any generalized eigenpair satisfies the limit eigenvalue equation on a suitable subgraph.
keywords:
$p$-Laplacian, variational eigenvalues, graph spectral theory, nonlinear eigenvalue problem, packing radii.
{MSCcodes}
35P30, 47J10, 05C12, 05C50
1 Introduction
We consider the problem of finding the eigenpairs of the $\infty$-Laplacian operator on graphs. While the continuous version of this problem has been studied extensively in the past few decades [31, 30, 35, 44, 22, 5], its discrete counterpart on graphs remains largely unexplored. In recent years, eigenpairs of the graph $\infty$-Laplace operator have found potential application in a variety of fields, including machine learning and data analysis [20], as well as $L^{1}$ optimal transport [4, 23] and the related problem of shape optimization [3], spectral clustering and semi-supervised learning [41, 18], and image manipulation [21, 20]. In the continuous setting, it is well known that the $p$-eigenpairs, $p\in[1,\infty]$, encode topological information about the domain [33, 37, 31, 30]. Moreover, in [10, 7], it is shown that different variational filtering methods produce a decomposition of the signal in terms of nonlinear eigenfunctions of the filter. In particular, in the finite dimensional case any such suitable filter can be written as $\left\|Ax\right\|_{\infty}$ for some proper matrix $A$. While this establishes a relationship with the discrete graph $\infty$-Laplacian eigenproblem, the study of these properties in the discrete setting is still largely unexplored.
The primary challenge of the $\infty$-Laplacian eigenvalue problem arises from the non-differentiability of the $\infty$-norm, which necessitates either the extension of traditional methods to the non-smooth regime or the development of entirely new approaches. When $1<p<\infty$, $p$-Laplacian eigenpairs are typically defined as critical points and values of the Rayleigh quotient $\mathcal{R}_{p}(f)=\|\nabla f\|_{p}/\|f\|_{p}$. However, a straightforward generalization of the $\infty$-eigenpairs using an $\infty$-Rayleigh quotient is complicated by the lack of smoothness and continuity inherent in the $\infty$-norm. Actually, we show in this work that different generalizations of the $p$-eigenpairs to the limit case $p=\infty$, lead to different notions of $\infty$-eigenpairs, each with distinct properties and implications. For example, in the continuous setting, the study of the $\infty$-eigenpairs, started in [31, 30], is based on the study of solutions of the limit $p$-Laplacian eigenvalue equation when $p$ goes to $\infty$. However, in the discrete setting, the formulation based on the eigenvalue equation is known to be different from the one based on the idea of a generalized critical point theory for nonsmooth functionals employed in [27, 11] for the graph $1$-Laplacian eigenvalue problem. Moreover, the generalized critical point theory has been successfully applied in recent works [9, 8] to investigate minimizers of $\|\cdot\|_{\infty}$ in both the $L^{\infty}$ and the $L^{2}$ settings, suggesting a promising direction for further exploration.
In this paper, we focus on the graph setting and examine the $\infty$-spectral problem from two complementary perspectives: the limit approach and the critical point theory. For clarity, we refer to the former as the $\infty$-limit eigenpairs and the latter as generalized $\infty$-eigenpairs.
Within the $\infty$-limit perspective, we address the extension to the discrete case of the results mainly obtained by Juutinen and co-authors [36, 31, 30, 35, 34]. More precisely, we are able to show that any $\infty$-limit eigenpair satisfies an $\infty$-limit eigenvalue equation. In addition, we prove that the $\infty$-limit of the $k$-th variational $p$-Laplacian eigenvalue can be bounded in terms of the $k$-th packing radius of the graph, i.e., the maximal radius that allows to inscribe $k$ distinct balls in the graph [26]. More general, we prove that the number of nodal domains of an $\infty$-eigenfunction can be used to bound the corresponding eigenvalue in terms of packing radii. Such results allow us to derive bounds from above and from below to the $\infty$-limit variational $p$-Laplacian eigenvalues in terms of the packing radii and, only for the first two $\infty$-limit variational eigenvalues, to establish equality with the first two packing radii.
Concerning the second perspective, we study the generalized $\infty$-eigenpairs and the variational $\infty$-eigenvalue, which are a subset of the former. We provide a comparison between the two formulations of $\infty$-eigenpairs in the graph setting. We prove, first of all, that the variational generalized $\infty$-eigenpairs satisfy the same approximation properties shared also by the $\infty$-limit variational eigenpairs [31, 30]. Then, we show that the $\infty$-limit eigenvalue problem is “stronger” than the generalized eigenvalue problem in the sense of the following inclusions:
(1) $\left\{\begin{array}[]{cc}&\text{Limit of }\\ &\Delta_{p}\text{-eigenpairs}\end{array}\right\}\subsetneq\left\{\begin{array}[]{cc}&\text{Solutions of the}\\ &\text{limit eigenvalue eq.}\end{array}\right\}\subsetneq\left\{\begin{array}[]{cc}&\text{Generalized }\\ &\infty\text{-eigenpairs }\end{array}\right\}\,.$
In more details, we consider the $\infty$-limit of $p$-Laplacian eigenpairs and prove that any accumulation point $(f,\Lambda)$ solves the system of equations
(2) $0=\begin{cases}\min\{\left\|\nabla_{\omega}f\right\|_{\infty}-\Lambda f(u)\;,\;\Delta_{\infty}f(u)\}\quad&\mathrm{if}\;f(u)>0\\ \Delta_{\infty}f(u)=0\quad&\mathrm{if}\;f(u)=0\\ \max\{-\left\|\nabla_{\omega}f\right\|_{\infty}-\Lambda f(u)\;,\;\Delta_{\infty}f(u)\}\quad&\mathrm{if}\;f(u)<0\end{cases}\,,$
but the inverse is not generally true. We are also able to prove that, if $(f_{k},\Lambda_{k})$ is an accumulation point of the $p$-sequence of the $k$-th variational eigenpairs of the $p$-Laplacian, then
$\Lambda_{k}\leq\frac{1}{R_{k}}\,,$
where $R_{k}$ is the $k$-th (sphere) packing radius of the graph, i.e., the maximal radius that allows to inscribe in the graph $k$ distinct balls of radius $R_{k}$. Moreover, denoting by $\mathcal{N}(f_{k})$ the number of nodal domains induced by $f_{k}$, we can prove that
$\frac{1}{R_{\mathcal{N}(f_{k})}}\leq\Lambda_{k}\,,$
In particular, since we prove that for $k=1,2$ the number of nodal domains exactly equals the index, i.e. $\mathcal{N}(f_{k})=k$, we are able to prove the following equalities:
$\Lambda_{k}=\frac{1}{R_{k}}\quad\mbox{if}\quad k=1,2\,.$
We obtain as a consequence that, as in the $p$-Laplacian case with $p\in(1,\infty)$ [43], the number of nodal domains induced by the $\infty$-eigenfunctions can be used to establish relationships between the corresponding eigenvalues and the packing radii.
In Section 4 we consider the generalized eigenvalue problem
(3) $0\in\partial\left\|\nabla_{\omega}f\right\|_{\infty}\cap\Lambda\partial\left\|f\right\|_{\infty}\,,$
where $\partial\left\|\nabla_{\omega}f\right\|_{\infty}$ and $\partial\left\|f\right\|_{\infty}$ denote the subdifferentials of the the two associated functions of $f$. Using the generalized Lusternik-Schnirelman theory [13], we define the Krasnoselskii variational $\infty$-eigenvalues $\Lambda_{k}^{\mathcal{K}}$ and prove that these eigenvalues satisfy:
$\Lambda_{k}^{\mathcal{K}}\leq\frac{1}{R_{k}}\quad\mbox{with}\quad\Lambda_{k}^{\mathcal{K}}=\frac{1}{R_{k}}\;\mathrm{if}\;k=1,2\,.$
Finally, we compare the $\infty$-limit and the generalized formulations and, using a geometrical characterization of the eigenvalues and eigenfunctions, we prove that the former is stronger than the latter in the sense defined above in (1). In addition, we are able to prove that any solution of the second formulation, up to considering a subgraph, solves the $\infty$-limit equation (2).
2 Notation and Preliminaries
In this section, we introduce our basic definitions and recall some preliminary results. In particular, since our goal is to prove discrete analogues of main results available for the continuous $\infty$-Laplacian eigenvalue problem, in this section, we also aim at establishing a solid framework for the discrete setting with definitions that find counterparts in the continuous case [43, 6, 1, 2].
2.1 Graph setting and $p$-Laplacian operators
We identify a graph $\mathcal{G}$ with a triplet $\mathcal{G}=(V,E,\omega)$, where $V$ denotes the set of nodes, $E\subset V\times V$ the set of edges, and $\omega:E\rightarrow\mathbb{R}$ is a weight on the edges. We identify an edge $(u,v)$ by the two nodes $u,v\in V$ that form its endpoints. In particular, the notation $u\sim v$ means that $(u,v)\in E$. A path $\Gamma$ connecting two nodes $v_{0},v_{n}\in V$ on the graph is a set of edges that connect $v_{0}$ with $v_{n}$, i.e., $\Gamma=\{v_{0}\sim v_{1}\sim\dots\sim v_{n}\}$. We say that a graph is connected if given any two nodes $u,v\in V$, there exists a path that joins them. Since we are only interested in undirected graphs, throughout the paper we assume that whenever $(u,v)\in E$ then also $(v,u)\in E$ and $\omega(u,v)=\omega(v,u)$. To simplify the notation, in the following we write $\omega_{uv}:=\omega(u,v)$. Moreover, we assume that $\omega_{uv}$ can be interpreted as the inverse of the edge length leading to the following notion of edge length, and consequently, distance between two nodes.
Definition 2.1 (Length of a path and distance between two nodes).
The length of a path $\Gamma$ is given by the sum of the reciprocals of the weights of the edges forming $\Gamma$,
$\mathrm{length}(\Gamma)=\sum_{i=1}^{n}\frac{1}{\omega_{v_{i-1}v_{i}}}\,.$
The distance between two nodes $u,v\in V$ is defined as the length of the shortest path connecting $u$ and $v$,
(4) $d(u,v)=\min\Big{\{}\sum_{i=1}^{n}\omega_{v_{i-1}v_{i}}^{-1}\,:n\in\mathbb{N},\;u=v_{0}\sim\dots\sim v_{n}=v\Big{\}}\,.$
Throughout the manuscript, if not otherwise specified, we use capital letters (latin or greek) to denote edge functions and lowercase letters to denote node functions.
We proceed now by introducing discrete analogues of the continuous differential operators. The gradient of a function $f$ defined on the nodes, $f:V\rightarrow\mathbb{R}$, is the edge function $\nabla_{\omega}f(u,v)$ given by
(5) $\nabla_{\omega}f(u,v)=\omega_{uv}\big{(}f(v)-f(u)\big{)}\,,$
for $(u,v)\in E$. Note that the linear operator $\nabla_{\omega}$ is represented by an $|E|\times|V|$ matrix called the oriented incidence matrix of the graph. Moreover, the assumption that the graph is undirected implies $\nabla_{\omega}f(u,v)=-\nabla_{\omega}f(v,u)$ for any $(u,v)\in E$. The set of the gradients of $f$ outgoing from node $u$, is called the local gradient $\left\{\nabla_{\omega}f\right\}(u)$ of $f$ at $u$:
$\left\{\nabla_{\omega}f\right\}(u):=\{\nabla_{\omega}f(u,v)\;|\;v\sim u\}\,.$
Note that, differently from the gradient operator in the continuous case, the cardinality of the local gradient is not fixed but depends on the degree of the node. Next, given an edge function $G:E\rightarrow\mathbb{R}$, we introduce its divergence as the node function defined by the following expression:
(6) ${\operatorname{div}G}(u)=-\frac{1}{2}\nabla_{\omega}^{T}G(u)=\frac{1}{2}\sum_{v\sim u}\omega_{uv}\left(G(u,v)-G(v,u)\right)$
Note that the above defined divergence operator satisfies the following “integration by parts” formula with respect to the scalar products $\left\langle\cdot,\cdot\right\rangle_{E}$ and $\left\langle\cdot,\cdot\right\rangle_{V}$ defined on the space of edge and node functions:
(7) $\left\langle G,\nabla_{\omega}f\right\rangle_{E}:=\frac{1}{2}\sum_{(u,v)\in E}G(u,v)\nabla_{\omega}f(u,v)=\sum_{u\in V}-{\operatorname{div}G}(u)f(u)=:\left\langle-{\operatorname{div}G},f\right\rangle_{V}\;.$
Observe that from the definition of the corresponding scalar products defined in (7), we can define the $p$-norms on the edge and node spaces as:
(8) $\left\|G\right\|_{p,E}^{p}=\frac{1}{2}\sum_{(u,v)\in E}|G(u,v)|^{p}\qquad\left\|f\right\|_{p,V}^{p}=\sum_{u\in V}|f(u)|^{p}\;,$
where we consider a one-half factor in edge-functions norms to take into account the fact that every edge has two possible directions and so it is counted twice. Finally, borrowing from [43, 16, 6], we define the $p$-Laplace operator as follows:
Definition 2.2 (Discrete $p$-Laplace Operator).
The discrete $p$-Laplace operator is given by:
$\Delta_{p}f(u):=-\operatorname{div}(|\nabla_{\omega}f|^{p-2}\odot\nabla_{\omega}f)=\sum_{v\sim u}\omega_{uv}^{p}|f(u)-f(v)|^{p-2}\left(f(u)-f(v)\right)\,,$
where $v\sim u$ means that $(u,v)\in E$, the absolute value in $|\nabla_{\omega}f|$ is considered entry-wise, and “ $\odot$” denotes the Hadamard or entry-wise product.
Next, we introduce the discrete counterpart of the homogeneous Dirichlet boundary conditions. To this aim, we designate an arbitrary subset of the nodes $B\subset V$ as the boundary of $\mathcal{G}$ and consider the set of node functions that are zero on $B$: $\mathcal{H}_{0}(V):=\{f\,|\;f(u)=0\;\forall u\in B\}$. From the study of the critical points of the $p$-Rayleigh quotient $\mathcal{R}_{p}:\mathcal{H}_{0}(V)\rightarrow\mathbb{R}$ given by
(9) $\mathcal{R}_{p}(f)=\frac{\left\|\nabla_{\omega}f\right\|_{p}}{\left\|f\right\|_{p}}$
we can write the discrete equivalent of the $p$-Laplacian eigenvalue equation with homogeneous Dirichlet boundary conditions [28, 38, 39] as:
(10) $\begin{cases}\Delta_{p}f(u)=\lambda|f|^{p-2}(u)f(u)\quad&u\in V_{\text{I}}\\ f(u)=0&u\in B\end{cases}\,.$
If $B\neq\emptyset$, we can introduce the distance from the boundary given by:
(11) $d_{B}(u)=\min_{v\in B}d(u,v)\,.$
In particular, if a graph has nonempty boundary we denote by $V_{\text{I}}:=V\setminus B$ the set of internal nodes and we say that the graph is connected if the subgraph induced by the internal nodes is connected. If not otherwise stated and independently of the presence of the boundary, throughout the paper we assume a graph to be connected. We define the $p$-Laplacian variational eigenvalues $\lambda_{k}(\Delta_{p})$ as in [43, 42] and write:
$\lambda_{k}(\Delta_{p})=\min_{A\in\mathcal{F}_{k}(S_{p})}\max_{f\in A}\mathcal{R}_{p}(f)\,,$
where $S_{p}:=\{f\in\mathcal{H}_{0}(V)|\;\|f\|_{p}=1\}$ and $\mathcal{F}_{k}(S_{p})$ is the set of the closed and symmetric subsets of $S_{p}$ with Krasnoselskii genus greater than or equal to $k$. We recall that the Krasnoselskii genus of a closed symmetric set $A$ is defined as follows [42, 25, 24]:
Definition 2.3 (Krasnoselskii Genus).
The Krasnoselskii Genus $\gamma(A)$ of a closed symmetric set $A$ is:
(12) $\gamma(A):=\begin{cases}0&\text{if $A=\emptyset$}\\ \inf\left\{h\in\mathbb{N}\,:\,\exists\,\varphi\in C(A,\mathbb{R}^{h}\setminus\{0\})\text{ s.t. }\varphi(x)=-\varphi(-x)\right\}&\\ \infty\quad&\text{if $\nexists\;h$ as above}\end{cases}.$
Given the $\infty$-norms of a node function $f$ and of its gradient:
(13) $\left\|f\right\|_{\infty}=\max_{u\in V}\{|f(u)|\}\,,\qquad\left\|\nabla_{\omega}f\right\|_{\infty}=\max_{(u,v)\in E}\{|\nabla_{\omega}f(u,v)|\},$
we can introduce the maximal node and edge sets:
$V_{\text{max}}(f)$ =$=\{u\in V\>:\>|f(u)|=\left\|f\right\|_{\infty}\}\,,$
$E_{\text{max}}(f)$ =$=\{(u,v)\in E\>:\>|\nabla_{\omega}f(u,v)|=\left\|\nabla_{\omega}f\right\|_{\infty}\}\,.$
Before considering the $\infty$-eigenvalue problem we need to introduce the corresponding $\infty$-Laplacian operator. We start by recalling the definition in the continuous setting [36] (differently from  [36] we are adding a minus sign to be consistent with our definition of the $p$-Laplacian operator that is positive definite differently from the one considered in [36]):
Definition 2.4 (Continuous $\infty$-Laplacian).
$\Delta_{\infty}f:=-\sum_{i,j}^{n}\frac{\partial f}{\partial x_{i}}\frac{\partial f}{\partial x_{j}}\frac{\partial^{2}f}{\partial x_{i}\partial x_{j}}=-(\nabla f)^{T}H(f)\nabla f\,,$
where $H(f)$ denotes the Hessian matrix of a twice differentiable function $f$ and $\nabla f$ its standard gradient.
Clearly, differently from the $p$-Laplacian operator with $p<\infty$, this definition cannot be extended to the discrete setting in a straight-forward manner because of the presence of the Hessian matrix. Nevertheless, a definition of the discrete $\infty$-Laplacian operator is given in [20, 21, 19] as follows:
Definition 2.5 (Discrete $\infty$-Laplacian Operator).
The discrete $\infty$-Laplacian operator can be written as:
$\Delta_{\infty}f(u)=\Big{(}\|\left\{\nabla_{\omega}f\right\}^{-}(u)\|_{\infty}-\|\left\{\nabla_{\omega}f\right\}^{+}(u)\|_{\infty}\Big{)}\,.$
This definition is consistent with the continous setting. Indeed, indicating with $f$ the limit as $p$ goes to $\infty$ of the solutions of $\Delta_{p}f_{p}=0$ with prescribed values on the boundary, it is easy to see that $f$ must satisfy $\Delta_{\infty}f=0$ for all $u\in V_{\text{I}}$ [20, 21, 19, 36]. In particular, the above definition is a direct consequence of the following reformulation of the discrete $p$-Laplacian operator [19]:
(14) $\Delta_{p}f(u):=$ $\sum_{v\in V}\omega_{uv}^{p}|f(u)-f(v)|^{p-2}\big{(}f(u)-f(v)\big{)}$
$=$ $\Big{(}\sum_{v\in V}\omega_{uv}^{p}\big{|}\big{(}f(u)-f(v)\big{)}^{+}\big{|}^{p-1}-\sum_{v\in V}\omega_{uv}^{p}\big{|}\big{(}f(u)-f(v)\big{)}^{-}\big{|}^{p-1}\Big{)}$
$=$ ($\Big{(}\|\left\{\nabla_{\omega^{\prime}}f\right\}^{-}(u)\|_{{p-1}}^{p-1}-\|\left\{\nabla_{\omega^{\prime}}f\right\}^{+}(u)\|_{{p-1}}^{p-1}\Big{)}\,,$
where $x^{+}:=\max\{x,0\}$, $x^{-}:=\max\{-x,0\}$ and $\omega^{\prime}_{uv}=\omega_{uv}^{\frac{p}{p-1}}$.
Remark 2.6.
Note that, given a graph with boundary $B$, using the inclusion operator $P:\mathcal{H}(V_{\text{I}})\rightarrow\mathcal{H}_{0}(V)$, it is easy to see that the space $\mathcal{H}_{0}(V)$ is isomorphic to the space $\mathcal{H}(V_{\text{I}}):=\{f:V_{\text{I}}\rightarrow\mathbb{R}\}$:
$Pf(u)=\begin{cases}f(u)\quad\forall u\in V_{\text{I}}\\ 0\quad\forall u\in B\end{cases}.$
This implies that $Pf$ is a $p$-Laplacian eigenfunction if $f$ satisfies:
$\frac{1}{2}P^{T}\nabla_{\omega}^{T}(|\nabla_{\omega}Pf|^{p-2}\odot\nabla_{\omega}Pf)(u)=\lambda|f(u)|^{p-2}f(u)\qquad\forall u\in V_{\text{I}},$
allowing us to forget the homogeneous Dirichlet boundary conditions. Thus, for the sake of notation, we will often refer to the eigenvalue problem in $\mathcal{H}(V_{\text{I}})$ and, with a small abuse of notation we will write $\nabla_{\omega}:\mathcal{H}(V_{\text{I}})\rightarrow\mathcal{H}(E)$, $\operatorname{div}:\mathcal{H}(E)\rightarrow\mathcal{H}(V_{\text{I}})$ and $\Delta_{p}:\mathcal{H}(V_{\text{I}})\rightarrow\mathcal{H}(V_{\text{I}})$ to denote
$\nabla_{\omega}:=\nabla_{\omega}P,\qquad\operatorname{div}:=P^{T}\operatorname{div},\qquad\Delta_{p}f:=-P^{T}\operatorname{div}(|\nabla_{\omega}Pf|^{p-2}\odot\nabla_{\omega}Pf)\,,$
where, in the last equalities, the $\nabla_{\omega}$ and $\operatorname{div}$ operators appearing in right hand sides are defined according to the definitions given in eqs. (5) and (6). Similarly, given a function $f\in\mathcal{H}(V_{\text{I}})$, we can modify the definition of the local gradient, of the $\infty$-Laplacian, and of the $p$-norm in accordance with the definition of the $p$-Rayleigh quotient:
$\nabla_{\omega}f(u):=\{\nabla_{\omega}Pf(u,v)|v\sim u\}\qquad u\in V_{\text{I}}$
$\Delta_{\infty}f(u)=\Big{(}\|\left\{\nabla_{\omega}Pf\right\}^{-}(u)\|_{\infty}-\|\left\{\nabla_{\omega}Pf\right\}^{+}(u)\|_{\infty}\Big{)}\qquad u\in V_{\text{I}}$
$\left\|f\right\|_{p,V}^{p}=\sum_{u\in V}|Pf(u)|^{p}=\sum_{u\in V_{\text{I}}}|f(u)|^{p}\qquad\left\|f\right\|_{\infty}=\max_{u\in V_{\text{I}}}\{|f(u)|\}.$
3 The $\infty$-limit eigenvalue problem
In this section we start discussing the $\infty$-Laplacian eigenvalue problem. In particular we concentrate on the $\infty$-limit eigenvalue problem”, i.e. we consider the limit as $p\rightarrow\infty$ of the $p$-Laplacian eigenvalue equation and we discuss its solutions. This approach has been thoroughly investigated in the continuous setting [31, 30, 35, 34, 22, 44]. Here, we start by recalling the main results in the continuous setting and then go on to prove analogous results in the discrete setting. We point out that, even if the results in the two settings are equivalent, the approaches to obtain them are different. We start by observing that, in both continuous and discrete settings, the limit of appropriate subsequences of $p$-Laplacian eigenpairs $(f(\Delta_{p}),\lambda(\Delta_{p})$ leads to an $\infty$-eigenfunction whose corresponding eigenvalue can be expressed formally as:
(15) $\Lambda=``\lim_{p\rightarrow\infty}\big{(}\lambda(\Delta_{p})\big{)}^{\frac{1}{p}}\,"\,.$
Moreover, we know that the set $\{\lambda^{\frac{1}{p}}(\Delta_{p})\}_{p}$ is bounded [16] and hence we can extract from it at least one convergent subsequence. Nevertheless there are no results about the convergence of entire sequences of $p$-Laplacian eigenvalues, neither when $\lambda(\Delta_{p})=\lambda_{k}(\Delta_{p})$ for a fixed value of $k$. In this last particular case we write $\Lambda_{k}$ to denote any accumulation point of the sequence $\{\lambda_{k}^{\frac{1}{p}}(\Delta_{p})\}_{p}$. In the continuous setting, we can write the $\infty$-eigenvalue equation for a general eigenvalue $\Lambda$ as [30]:
(16) $0=\begin{cases}\min\{|\nabla_{\omega}f|-\Lambda f,\Delta_{\infty}f\}\quad&\mathrm{if}\;f>0\\ \Delta_{\infty}f&\mathrm{if}\;f=0\\ \max\{-|\nabla_{\omega}f|-\Lambda f,\Delta_{\infty}f\}\quad&\mathrm{if}\;f<0\end{cases}\,,$
where the above equation has to be understood in the sense of viscosity solutions (see [31] for a short and clear discussion about viscosity solutions in this setting). In particular it is possible to show that if $(f,\Lambda)=\lim_{p\rightarrow\infty}(f_{p},\lambda_{p}^{\frac{1}{p}})$, where $(f_{p},\lambda_{p})$ is an eigenpair of the $p$-Laplacian for a given $p$ then $(f,\Lambda)$ solves (16). In addition, whenever a pair $(f,\Lambda)$ solves eq.(16), then
$\Lambda=\frac{\|\nabla_{\omega}f\|_{\infty}}{\|f\|_{\infty}}\;.$
By means of this equation it is possible to relate the limit of the variational $p$-Laplacian eigenvalues to the packing radii of $\Omega$. Indeed, given a bounded domain $\Omega\in\mathbb{R}^{n}$ and an integer $k$, the $k$-th packing radius of $\Omega$ is defined as the maximal radius $r\in\mathbb{R}$ that allows inscribing $k$ distinct balls in $\Omega$, i.e.:
$R_{k}:=\sup\{r\in\mathbb{R}\text{ s.t. }\exists B_{r}(x_{1}),\dots,B_{r}(x_{k})\subset\Omega,\;B_{r}(x_{i})\cap B_{r}(x_{j})=\emptyset\;\forall i,j=1,\dots,k\}\,,$
where $B_{r}(x_{i})$ is the $n$-dimensional ball of radius $r$ centered in $x_{i}$, $x_{i}\in\Omega$. Then, given $\Lambda_{k}$ an accumulation point of the sequence of the $k$-th variational eigenvalues of the $p$-Laplacian, i.e.
$\Lambda_{k}=\lim_{h\rightarrow\infty}\lambda_{k}^{\frac{1}{p_{h}}}(\Delta_{p_{h}})\,,$
it is possible to relate such value to $R_{k}$ (see [30, 31]) and obtain the following inequality:
$\Lambda_{k}\leq\frac{1}{R_{k}}\,.$
For the particular cases $k=1,2$, also the opposite inequality holds and the sequences $\{\lambda_{1,2}^{\frac{1}{p}}(\Delta_{p})\}_{p=2}^{\infty}$ converge. Thus, we can write:
$\Lambda_{1}=\frac{1}{\|d_{\partial\Omega}\|_{\infty}}\,,\qquad\Lambda_{2}=\frac{1}{R_{2}}\,,$
where $\partial\Omega$ denotes the boundary of the domain $\Omega$. Several studies have concerned the $\infty$-eigenfunctions associated to the first eigenvalue $\Lambda_{1}$ (see [36, 44, 32, 29]). Indeed, despite the fact that the boundary distance function $d_{\partial\Omega}$ is always a minimizer of $\mathcal{R}_{\infty}$, it is not always an eigenfunction, meaning that, depending on $\Omega$, the pair $(d_{\partial\Omega},1/\|d_{\partial\Omega}\|_{\infty})$ may not solve eq. (16). In addition, the first eigenfunction may not be unique and there may be first eigenfunctions that solve the $\infty$-limit eigenvalue equation without beeing limits of eigenfunctions of the $p$-Laplacian.
3.1 The $\infty$-limit eigenvalue equation on a graph
In this section we formulate discrete analogues of the $\infty$-eigenvalue eq. (16) and discuss their solutions. Note that, given a a function $f\in\mathcal{H}(V_{\text{I}})$ , using the equalities in (14) we can rewrite the eigenvalue equation of the $p$-Laplacian given in (10) as:
(17a) {($\begin{cases}\left(\left\|\left\{\nabla_{\omega^{\prime}}f\right\}^{-}(u)\right\|_{p-1}^{p-1}-\left\|\left\{\nabla_{\omega^{\prime}}f\right\}^{+}(u)\right\|_{p-1}^{p-1}\right)^{\frac{1}{p-1}}=\lambda^{\frac{1}{p-1}}f(u)\\[4.73611pt] \left\|\left\{\nabla_{\omega^{\prime}}f\right\}^{-}(u)\right\|_{p-1}-\left\|\left\{\nabla_{\omega^{\prime}}f\right\}^{+}(u)\right\|_{p-1}>0\end{cases}\!\!\!\!\!\!\!\!\quad\text{if }f(u)>0$
(17b) {($\begin{cases}\left(\left\|\left\{\nabla_{\omega^{\prime}}f\right\}^{+}(u)\right\|_{p-1}^{p-1}-\left\|\left\{\nabla_{\omega^{\prime}}f\right\}^{-}(u)\right\|_{p-1}^{p-1}\right)^{\frac{1}{p-1}}=-\lambda^{\frac{1}{p-1}}f(u)\\[4.73611pt] \left\|\left\{\nabla_{\omega^{\prime}}f\right\}^{-}(u)\right\|_{p-1}-\left\|\left\{\nabla_{\omega^{\prime}}f\right\}^{+}(u)\right\|_{p-1}^{p-1}<0\end{cases}\!\!\!\!\!\!\!\!\quad\text{if }f(u)<0$
(17c) $\phantom{.}\,\>\left\|\left\{\nabla_{\omega^{\prime}}f\right\}^{-}(u)\right\|_{p-1}-\left\|\left\{\nabla_{\omega^{\prime}}f\right\}^{+}(u)\right\|_{p-1}=0\,\,\qquad\text{if }f(u)=0$
where, as in eq. (14), $\omega^{\prime}_{uv}:=\omega_{uv}^{\frac{p}{p-1}}$. In the following Theorem we prove that any accumulation point of a sequence of $p$-Laplacian eigenpairs satisfies the $\infty$-limit eigenvalue equation. Before stating the Theorem, we would like to observe that, as shown in [16], any $p$-Laplacian eigenvalue $\lambda(\Delta_{p})$ satisfies the following inequality:
$\lambda(\Delta_{p})\leq\max_{u\in V_{\text{I}}}2^{p-1}\sum_{v\sim u}\omega_{uv}^{p}\,.$
Thus, the set of the $p$-roots of the $p$-Laplacian eigenvalues as $p$ varies is bounded, i.e.:
there exists a constant $\text{there exists a constant }C>0\text{ s.t. }\lambda^{\frac{1}{p}}<C\text{ for all }\lambda\in\sigma(\Delta_{p})\,,$
where $\sigma(\Delta_{p})$ denotes the spectrum of the graph $p$-Laplacian.
Theorem 3.1.
Let $(f_{p_{j}},\lambda_{{p_{j}}})$ be a sequence of $p$-Laplacian eigenpairs and assume that $\lim_{j\rightarrow\infty}(f_{p_{j}},\lambda_{p_{j}}^{\frac{1}{p_{j}}})=(f,\Lambda)$. Then the pair $(f,\Lambda)$ satisfies the following set of equations:
(18) $0=\begin{cases}\min\{\|\left\{\nabla_{\omega}f\right\}(u)\|_{\infty}-\Lambda f(u)\,,\,\Delta_{\infty}f(u)\}\quad&\text{if }f(u)>0\\ \Delta_{\infty}f(u)\quad&\text{if }f(u)=0\\ \max\{-\|\left\{\nabla_{\omega}f\right\}(u)\|_{\infty}-\Lambda f(u)\,,\,\Delta_{\infty}f(u)\}\quad&\text{if }f(u)<0\end{cases}\,.$
Proof 3.2.
To prove the assertion, consider the case $f(u)>0$. Then, for any $j$ large enough, by (17a), we can write
(19) $\begin{cases}\displaystyle{\|\left\{\nabla_{\omega^{\prime}}f_{p_{j}}\right\}^{-}(u)\|_{p_{j}-1}\Big{(}1-\Big{(}\frac{\|\left\{\nabla_{\omega^{\prime}}f_{p_{j}}\right\}^{+}(u)\|_{p_{j}-1}}{\|\left\{\nabla_{\omega^{\prime}}f_{p_{j}}\right\}^{-}(u)\|_{p_{j}-1}}\Big{)}^{p_{j}-1}\Big{)}^{\frac{1}{p_{j}-1}}=\lambda_{p_{j}}^{\frac{1}{p_{j}-1}}f_{p_{j}}(u)\,,}\\ \|\left\{\nabla_{\omega^{\prime}}f_{p_{j}}\right\}^{-}(u)\|_{p_{j}-1}>\|\left\{\nabla_{\omega^{\prime}}f_{p_{j}}\right\}^{+}(u)\|_{p_{j}-1}\,.\end{cases}$
Taking the limit we have
(20) $\begin{cases}\displaystyle\|\left\{\nabla_{\omega}f\right\}(u)\|_{\infty}\lim_{j\rightarrow\infty}\Big{(}1-\Big{(}\frac{\|\left\{\nabla_{\omega^{\prime}}f_{p_{j}}\right\}^{+}(u)\|_{p_{j}-1}}{\|\left\{\nabla_{\omega^{\prime}}f_{p_{j}}\right\}^{-}(u)\|_{p_{j}-1}}\Big{)}^{p_{j}-1}\Big{)}^{\frac{1}{p_{j}-1}}=\Lambda f(u)\\ \|\left\{\nabla_{\omega}f\right\}^{-}(u)\|_{\infty}\geq\|\left\{\nabla_{\omega}f\right\}^{+}(u)\|_{\infty}\end{cases},$
where, we have used the equality $\|\left\{\nabla_{\omega}f\right\}^{-}(u)\|_{\infty}=\|\left\{\nabla_{\omega}f\right\}(u)\|_{\infty}$ that follows trivially from the second inequality in (19). From this same inequality, we observe that
$\lim_{j\rightarrow\infty}\bigg{(}1-\bigg{(}\frac{\|\left\{\nabla_{\omega^{\prime}}f_{p_{j}}\right\}^{+}(u)\|_{p_{j}-1}}{\|\left\{\nabla_{\omega^{\prime}}f_{p_{j}}\right\}^{-}(u)\|_{p_{j}-1}}\bigg{)}^{p_{j}-1}\bigg{)}^{\frac{1}{p_{j}-1}}\leq 1,$
and thus, if $\left\|\left\{\nabla_{\omega}f\right\}^{-}(u)\right\|_{\infty}\gneqq\|\left\{\nabla_{\omega}f\right\}^{+}(u)\|_{\infty}$, we obtain:
$\lim_{j\rightarrow\infty}\bigg{(}1-\bigg{(}\frac{\|\left\{\nabla_{\omega^{\prime}}f_{p_{j}}\right\}^{+}(u)\|_{p_{j}-1}}{\|\left\{\nabla_{\omega^{\prime}}f_{p_{j}}\right\}^{-}(u)\|_{p_{j}-1}}\bigg{)}^{p_{j}-1}\bigg{)}^{\frac{1}{p_{j}-1}}=1.$
Replacing the last equation in (20) we obtain:
$\|\left\{\nabla_{\omega}f\right\}(u)\|_{\infty}=\Lambda f(u)\,.$
Vice versa, if ${\displaystyle\lim_{j\rightarrow\infty}}\Big{(}1-\Big{(}\frac{\|\left\{\nabla_{\omega^{\prime}}f_{p_{j}}\right\}^{+}(u)\|_{p_{j}-1}}{\|\left\{\nabla_{\omega^{\prime}}f_{p_{j}}\right\}^{-}(u)\|_{p_{j}-1}}\Big{)}^{p_{j}-1}\Big{)}^{\frac{1}{p_{j}-1}}<1$, then necessarily we have
$\|\left\{\nabla_{\omega}f\right\}^{-}(u)\|_{\infty}=\|\left\{\nabla_{\omega}f\right\}^{+}(u)\|_{\infty}\,,$
showing that (18) is satisfied. The cases $f(u)=0$ and $f(u)<0$ can be proved analogously, concluding the proof.
We point out that eq. (18) is very similar to the continuous analogue in (16), with the only difference being the absolute value replaced by the $\infty$-norm of the local gradient. Next we show that whenever $(f,\Lambda)$ satisfies the limit eigenvalue eq .(18), $\Lambda$ matches the value of the $\infty$-Rayleigh quotient computed in $f$, i.e. $\mathcal{R}_{\infty}(f)$.
Proposition 3.3.
Assume that $(f,\Lambda)$ satifies eq. (18), $f$ is not a constant function, and $u\in V_{\text{max}}(f)$. Then $\|\nabla_{\omega}f\|_{\infty}=\|\left\{\nabla_{\omega}f\right\}(u)\|_{\infty}=\Lambda|f(u)|=\Lambda\|f\|_{\infty}$, i.e.
$\Lambda=\frac{\|\nabla_{\omega}f\|_{\infty}}{\|f\|_{\infty}}\,.$
Proof 3.4.
Assume $u\in V_{\text{I}}$ to be a node such that $|f(u)|=\|f\|_{\infty}$, i.e., $u\in V_{\text{max}}(f)$, and, without loss of generality, that $f(u)>0$. Then (18) ensures that:
$\|\left\{\nabla_{\omega}f\right\}(u)\|_{\infty}\geq\Lambda f(u)>0\,.$
Since $u$ is a maximum point, we have that $\|\left\{\nabla_{\omega}f\right\}^{+}(u)\|_{\infty}=0$, which, together with the previous inequality, yields:
$\Delta_{\infty}f(u)>0\,.$
Thus, from (18) we necessarily have:
$\|\left\{\nabla_{\omega}f\right\}(u)\|_{\infty}-\Lambda f(u)=0\,.$
The last equality allows us to observe that, if the thesis holds for some $w\in V_{\text{max}}(f)$, it holds as well for all the other $u\in V_{\text{max}}(f)$. Indeed
$\frac{\|\left\{\nabla_{\omega}f\right\}(u)\|_{\infty}}{|f(u)|}=\frac{\|\left\{\nabla_{\omega}f\right\}(u)\|_{\infty}}{\|f\|_{\infty}}=\Lambda=\frac{\|\left\{\nabla_{\omega}f\right\}(w)\|_{\infty}}{|f(w)|}=\frac{\|\nabla_{\omega}f\|_{\infty}}{\|f\|_{\infty}}\,.$
To prove the existence of such a node $w$, we proceed as follows. Let $u_{0}\in V_{\text{I}}$ be a node admitting an edge $(v_{0},u_{0})$ with $|\nabla_{\omega}f(v_{0},u_{0})|=\|\nabla_{\omega}f\|_{\infty}=\|\left\{\nabla_{\omega}f\right\}(u_{0})\|_{\infty}=\|\left\{\nabla_{\omega}f\right\}(v_{0})\|_{\infty}$. Then, if $\Delta_{\infty}f(u_{0})\neq 0$, we can write:
$\Lambda=\frac{\|\left\{\nabla_{\omega}f\right\}(u_{0})\|_{\infty}}{|f(u_{0})|}=\frac{\|\nabla_{\omega}f\|_{\infty}}{|f(u_{0})|}\geq\frac{\|\left\{\nabla_{\omega}f\right\}(u)\|_{\infty}}{|f(u)|}=\frac{\|\left\{\nabla_{\omega}f\right\}(u)\|_{\infty}}{\|f\|_{\infty}}=\Lambda\,,$
where $u$ is any node satisfying $|f(u)|=\|f\|_{\infty}$, yielding in particular $|f(u_{0})|=\|f\|_{\infty}$ and thus the thesis. On the other hand, if $\Delta_{\infty}f(u_{0})=\Delta_{\infty}f(v_{0})=0$, assuming without loss of generality $f(u_{0})>f(v_{0})$ (indeed the role of $u_{0}$ and $v_{0}$ is interchangeable), there exists $u_{1}\sim u_{0}$ such that $f(u_{1})>f(u_{0})$ with
$\omega_{u_{0}u_{1}}\big{(}f(u_{1})-f(u_{0})\big{)}$ =$=\|\left\{\nabla_{\omega}f\right\}^{-}(u_{0})\|_{\infty}$
=$=\|\left\{\nabla_{\omega}f\right\}^{+}(u_{0})\|_{\infty}=\omega_{u_{0},v_{0}}\big{(}f(u_{0})-f(v_{0})\big{)}=\|\nabla_{\omega}f\|_{\infty}\,.$
Thus there exists another edge $(u_{0},u_{1})$ such that $|\nabla_{\omega}f(u_{0},u_{1})|=\|\nabla_{\omega}f\|_{\infty}$ and $f(u_{1})>f(u_{0})>f(v_{0})$. Iterating this procedure, by the finiteness of the graph and the previous argument, there must exist a node $u_{k}$ such that:
$\|\left\{\nabla_{\omega}f\right\}(u_{k})\|_{\infty}=\|\nabla_{\omega}f\|_{\infty}=\Lambda|f(u_{k})|=\Lambda\|f\|_{\infty}\,.$
3.2 Topological Properties of the $\infty$-eigenpairs
[…]
Proposition 3.5.
Let $(f,\Lambda)$, be a pair satisfying (18) with $f$ not a constant function. Then for any $u\in V_{\text{max}}(f)$ there exists a path $\Gamma=\{(u_{i},u_{i+1})\}_{i=1}^{n-1}$ starting from $u$ ($u_{1}=u$) such that :
  1. 1.
    $f$ is monotone along $\Gamma$ and for all $11\leq k\leq n-$ 1, $\nabla_{\omega}f(u_{k},u_{k+1})=\|\nabla_{\omega}f\|_{\infty}$;
  2. 2.
    either $\begin{cases}u_{n}\in B\\ \displaystyle{\Lambda=\frac{1}{\mathrm{length}(\Gamma)}}\end{cases}$ or $\begin{cases}f(u_{n})=-f(u)\\ \displaystyle{\Lambda=\frac{2}{\mathrm{length}(\Gamma)}}\end{cases}$;
  3. 3.
    ${\Lambda=\min\bigg{\{}\frac{1}{d_{B}(u)},\min_{\{v\,|\,f(v)=-f(u)\}}\frac{2}{d(u,v)}\bigg{\}}}\,.$
Proof 3.6.
The first two points follow from Proposition 3.3 and eq. (18). Indeed, given $u_{1}=u$, by Proposition 3.3 there exists $u_{2}\sim u_{1}$ such that $|\nabla_{\omega}f(u_{1},u_{2})|=\|\nabla_{\omega}f\|_{\infty}$. Then, from (18) we conclude that one of the three following conditions must hold:
$f(u_{2})=-f(u_{1})\qquad\mbox{or}\qquad u_{2}\in B\qquad\mbox{or}\qquad|f(u_{2})|<|f(u_{1})|\,.$
In the last case, since $\|\nabla_{\omega}f(u_{2})\|_{\infty}=\|\nabla_{\omega}f\|_{\infty}$, necessarily $\Delta_{\infty}f(u_{2})=0$, i.e. there exists $u_{3}\sim u_{2}$ such that $|\nabla_{\omega}f(u_{2},u_{3})|=\|\nabla_{\omega}f\|_{\infty}$ and, assuming w.l.o.g. $f(u_{1})>f(u_{2})$, it follows $f(u_{2})>f(u_{3})$. Iterating this procedure, the finiteness of the graph ensures that a path $\Gamma$ such that $u_{n}\in B$ or $f(u_{n})=-f(u_{1})$ exists. The following equalities
$\frac{f(u_{i})-f(u_{i+1})}{\|\nabla_{\omega}f\|_{\infty}}=\frac{1}{\omega_{u_{i},u_{i+1}}}\,\>\forall i=1,\dots,n-1$
provide an explicit expression for the length of the path $\Gamma$:
(21) $\mathrm{length}(\Gamma)=\sum_{i}\frac{1}{\omega_{u_{i},u_{i+1}}}=\frac{f(u_{1})-f(u_{n})}{\|\nabla_{\omega}f\|_{\infty}}\,.$
This last equation, together with the equality $\Lambda=\|\nabla_{\omega}f\|_{\infty}/|f(u_{1})|$, yields $\Lambda=1/\mathrm{length}(\Gamma)$ if $u_{n}\in B$ and $\Lambda=2/\mathrm{length}(\Gamma)$ if $f(u_{n})=-f(u_{1})$
To conclude we are left to prove that i) if $u_{n}\in B$ then $\mathrm{length}(\Gamma)=d_{B}(u_{1})$ and ii) $\mathrm{length}(\Gamma)\leq 2d(u_{1},v)$ for all $v$ such that $f(v)=-f(u_{1})$, together with the analogous expressions for the case $f(u_{n})=-f(u_{1})$. To this aim, assume by contradiction that there exists a path $\Gamma^{\prime}=\{(v_{i},v_{i+1})\}_{i=1}^{m-1}$ with $v_{1}=u_{1}$, $v_{m}\in B$, and $\sum_{i=1}^{m-1}\omega_{v_{i}v_{i+1}}^{-1}<\sum_{i=1}^{n-1}\omega_{u_{i}u_{i+1}}^{-1}$ and observe that
$\omega_{v_{i}v_{i+1}}|f(v_{i})-f(v_{i+1})|=|\nabla_{\omega}f(v_{i},v_{i+1})|\leq\|\nabla_{\omega}f\|_{\infty},\quad\forall i=1,\dots,m-1\,,$
Then, using the triangular inequality and then equality (21), and the assumptions $u_{n},v_{m}\in B$ (i.e. $f(u_{n})=f(u_{m})=0$), we obtain the following contradiction:
$|f(u_{1})|\leq\sum_{i=1}^{m-1}|f(v_{i})-f(v_{i+1})|$ $\leq\|\nabla_{\omega}f\|_{\infty}\sum_{i=1}^{m-1}\omega_{v_{i}v_{i+1}}^{-1}$
$<\|\nabla_{\omega}f\|_{\infty}\sum_{i=1}^{n-1}\omega_{u_{i}u_{i+1}}^{-1}=|f(u_{1})|\,.$
The inequality $\mathrm{length}(\Gamma)\leq 2\,d(u_{1},v)$ for all $v$ such that $f(v)=-f(u_{1})$, can be proved with a similar argument, i.e., assuming the existence of a path $\Gamma^{\prime}$, from $u_{1}$ to $v$, with $f(v)=-f(u_{1})$, such that $\mathrm{length}(\Gamma^{\prime})<2\,\mathrm{length}(\Gamma)$. The final case $f(u_{n})=-f(u_{1})$ can be treated analogously.
We now turn our attention to the study of the accumulation points of sequences of $p$-Laplacian variational eigenvalues. We start by introducing some necessary preliminary results that are of independent interest. First of all we observe that the minimum of the $\infty$-Rayleigh quotient
(22) $\mathcal{R}_{\infty}(f)=\frac{\|\nabla_{\omega}f\|_{\infty}}{\|f\|_{\infty}}$
can be easily computed, as in the continuous case [30, 9], considering the distance from the boundary, $d_{B}$ given in (11). Before proving this result, let us consider the following trivial inequality. Assume the distance between the nodes $u$ and $v$ is realized along the path $\Gamma=\{(v_{i},v_{i+1})\}_{i=1}^{n-1}$, with $v_{1}=u$ and $v_{n}=v$, i.e., $d(u,v)=\sum_{i=1}^{m}1/\omega_{v_{i}v_{i+1}}\,.$ Then, the trianguar inequality yields the following Lipschitz condition:
(23) $|f(u)-f(v)|$ $\leq|\sum_{i=1}^{m}|f(v_{i})-f(v_{i+1})|\leq\|\nabla_{\omega}f\|_{\infty}\sum_{i=1}^{m}\frac{1}{\omega_{v_{i}v_{i+1}}}=\|\nabla_{\omega}f\|_{\infty}d(u,v)\,.$
The next result shows that the boundary distance function $d_{B}$ defined in (11) is a minimizer of the Rayleigh quotient.
Proposition 3.7.
The boundary distance function $d_{B}\in V$ realizes the minimunm of the $\infty$-Rayleigh quotient:
$d_{B}\in\operatorname*{arg\>min}_{f\in\mathcal{H}_{0}(V)}\frac{\|\nabla_{\omega}f\|_{\infty}}{\|f\|_{\infty}}\,,$
Proof 3.8.
First we prove that $\|\nabla_{\omega}d_{B}\|_{\infty}\leq 1$. Indeed, given a node $u\in V$ and an edge $(u,v)\in E$ connected to it, the gradient of the distance function $d_{B}$ on this edge can be estimated as:
$\nabla_{\omega}d_{B}(u,v)=\omega_{uv}\big{(}d_{B}(v)-d_{B}(u)\big{)}\leq\omega_{uv}(\omega_{uv})^{-1}=1\,.$
Thus, to conlude it is sufficient to show that the boundary distance function satisfies:
$d_{B}\in\operatorname*{arg\>max}_{f\in\mathcal{F}:\|\nabla_{\omega}f\|_{\infty}\leq 1}\|f\|_{\infty}\,.$
To this end, for any function $f$ with $f\in\mathcal{F}$ and $\|\nabla_{\omega}f\|_{\infty}\leq 1$ and for any node $u\in V$, let $v\in B$ be the boundary node such that $d(u,v)=d_{B}(u)$. Then, using (23) we can write:
$|f(u)|=|f(u)-f(v)|\leq d(u,v)=d_{B}(u)\,.$
The last equation clearly implies:
$d_{B}\in\operatorname*{arg\>max}_{f\in\mathcal{H}_{0}(V):\|\nabla_{\omega}f\>\|_{\infty}\leq 1}\|f\|_{\infty}.$
We can also show that the minimum value of the Rayleigh quotient is equal to the radius of the largest ball that can be inscribed in $\mathcal{G}$.
Corollary 3.9.
$\min_{f\in\mathcal{H}_{0}(V)}\frac{\|\nabla_{\omega}f\|_{\infty}}{\|f\|_{\infty}}=\frac{1}{R_{1}}\,,$
where $R_{1}$ is defined as $R_{1}=\sup_{u\in V_{\text{I}}}\{d_{B}(v)\}$.
Proof 3.10.
Because of Proposition 3.7, it is enough to compute $\|\nabla_{\omega}d_{B}\|_{\infty}$ and $\|d_{B}\|_{\infty}$. Trivially $\|d_{B}\|_{\infty}=R_{1}$. From the proof of Proposition 3.7, we already know that $\big{|}\nabla_{\omega}d_{B}(u,v)\big{|}\leq 1\;\>\forall(u,v)\in E$. Moreover, it is easy to prove (see e.g. [9]) that:
$|\nabla_{\omega}d_{B}(u,v)|=1\quad\Longleftrightarrow\quad v\in\Gamma(u,B)\,\>or\,u\in\Gamma(v,B)\,,$
where $\Gamma(u,B)$ is the shortest path from the node $u$ to the boundary. Thus, necessarily $\|\nabla_{\omega}d_{B}\|_{\infty}=1$.
The next theorem shows that, given $(f,\Lambda)$ solution to the $\infty$-limit eigenvalue eq. (18), the eigenvalue $\Lambda$ can be regarded as the maximal radius that allows one to inscribe as many balls in the graph as the number of nodal domains induced by $f$. Recall that the nodal domains induced by $f$ are the maximal connected subgraphs where $f$ is strictly positive or negative. The proof of the theorem relies on results related to the $p$-Laplacian taken from [43, 16], more specifically when $p=1$ as specialized in [12]
Theorem 3.11.
Let $(f,\Lambda)$ be an eignepair that satisfies eq. (18), and assume that $f$ induces $k$ nodal domains. Then there exist $v_{1},\dots,v_{k}\in V$ and $r_{k}>0$ such that
$d(v_{i},v_{j})\geq 2r_{k}\quad\quad d_{B}(v_{i})\geq r_{k}\quad\text{ for all }i,j=1,\dots,k,i\neq j\,.$
Moreover
$\Lambda=\frac{1}{r_{k}}\,.$
Proof 3.12.
Consider first the case of $f$ strictly positive. Then, given $v_{1}\in V_{\text{max}}(f)$, from Proposition 3.5 there exists a path $\Gamma=\{(u_{i},u_{i+1})\}_{i=1}^{n-1}$ starting from $u_{1}=v_{1}$ such that $u_{n}\in B$ and $\mathrm{length}(\Gamma)=d_{B}(v_{1})=\Lambda^{-1}$. Thus, the node $v_{1}$ and the radius $r_{1}=\Lambda^{-1}$ satisfy the thesis.
Consider now the case of a function $f$ satisfying (18) and assume $\{A_{i}\}$, $\{B_{j}\}$ to be its nodal domains in such a way that:
$f(u)>0\quad\forall u\in\cup_{i}A_{i}\,,\qquad f(u)<0\quad\forall u\in\cup_{j}B_{j}\,.$
Starting from the original graph $\mathcal{G}$ and the function $f$, we aim at defining a new disconnected graph $\mathcal{G}^{\prime}$ such that:
  • $\mathcal{G}^{\prime}$ has as many connected components, $\mathcal{G}_{h}$, as the nodal domains of $f$,
  • for any $h$, $(f|_{\mathcal{G}_{h}},\Lambda)$ satisfy the $\infty$-limit eigenvalue equation (18) on $\mathcal{G}_{h}$,
  • $f|_{\mathcal{G}_{h}}$ is strictly positive or strictly negative.
To this end, we consider all the nodes $u$ such that $f(u)=0$ as boundary nodes of the new graph, $u\in B(\mathcal{G}^{\prime})$. Furthermore, for any edge $(u,v)$ such that $f(u)f(v)<0$, we add a boundary node $w\in B(\mathcal{G}^{\prime})$ and replace the edge $(u,v)$ by two edges $(u,w)$ and $(w,v)$ with weights $\omega_{uw}=\omega_{uv}\big{(}1-f(v)/f(u)\big{)}$ and $\omega_{wv}=\omega_{uv}\big{(}1-f(u)/f(v)\big{)}$. Observe that, since $f(w)=0$ ($w\in B(\mathcal{G}^{\prime})$), the gradient on $(u,v)$ is the same as the gradients on $(u,w)$ and $(w,v)$, i.e.:
(24) $\nabla_{\omega}f(u,v)=\omega_{uv}\big{(}f(v)-f(u)\big{)}=\omega_{uw}\big{(}0-f(u)\big{)}=\nabla_{\omega}f(u,w),$
$\nabla_{\omega}f(v,u)=\omega_{uv}\big{(}f(u)-f(v)\big{)}=\omega_{vw}\big{(}0-f(v)\big{)}=\nabla_{\omega}f(v,w),$
and the weights are conjugate:
(25) $\frac{1}{\omega_{uw}}+\frac{1}{\omega_{vw}}=\frac{1}{\omega_{uv}}\,.$
Note that the internal nodes of the new graph $\mathcal{G}^{\prime}$ are the nodes of $\mathcal{G}$ where $f$ is non zero. In addition, $\mathcal{G}^{\prime}=(\cup_{h=1}^{n}\mathcal{G}_{h})\cup B(\mathcal{G}^{\prime})$ with $\big{(}\mathcal{G}_{h_{1}}\cap\mathcal{G}_{h_{2}}=\emptyset$ for any pair of distinct indices $h_{1}$ and $h_{2}$ varying between 1 and $n$. Moreover, every $\mathcal{G}_{h}$ matches one of the nodal domains induced by $f$, meaning that its internal nodes are all and only the nodes that belong to some nodal domain induced by $f$ on $\mathcal{G}$.
We now observe that for any internal node $u$ of $\mathcal{G}^{\prime}$, from (24), the local gradient, $\left\{\nabla_{\omega}f\right\}(u)$ is unchanged, implying that $(f|_{\mathcal{G}_{h}},\Lambda)$ satisfies the $\infty$-limit equation (18) with respect to the graph $\mathcal{G}_{h}$. Thus, since $f|_{\mathcal{G}_{h}}$ is strictly positive or negative, the first part of the proof guarantees that, for any $h$, there exists $v_{h}\in\mathcal{G}_{h}$ such that
$\Lambda=\frac{1}{d_{B(\mathcal{G}^{\prime})}(v_{h})}\,.$
We conclude the proof of the Theorem by observing that:
$\frac{2}{\Lambda}=d_{B(\mathcal{G}^{\prime})}(v_{h_{1}})+d_{B(\mathcal{G}^{\prime})}(v_{h_{2}})\leq d(v_{h_{1}},v_{h_{2}})\,.$
To prove this last assertion, we first recall that $\mathcal{G}_{h_{1}}\cap\mathcal{G}_{h_{2}}=\emptyset$ and that, by construction eq. (25) holds, i.e., we have not changed distances among internal nodes. Now, we let $\Gamma$ be the shortest path that joins $v_{h_{1}}$ and $v_{h_{2}}$ in $\mathcal{G}$, i.e. $d(v_{h_{1}},v_{h_{2}})=\mathrm{length}(\Gamma)$. Then, in $\mathcal{G}^{\prime}$, the path $\Gamma$ has necessarily been replaced by some path $\Gamma^{\prime}$ that crosses $B(\mathcal{G}^{\prime})$ and such that $\mathrm{length}(\Gamma^{\prime})=\mathrm{length}(\Gamma)$.
Finally, we conclude this section by studying the relationships between the limits of the $p$-Laplacian variational eigenvalues, $\{\Lambda_{k}\}_{k}$, and the packing radii of $\mathcal{G}$ [26], $\{R_{k}\}_{k}$, i.e. the maximal radii that allow one to inscribe a prescribed number of disjoint balls in the graph. We start by defining the $k$-th packing radius of the graph $\mathcal{G}$ as follows.
Definition 3.13 ($k$-th packing radius $R_{k}$).
$R_{k}:=\max\left\{r>0:\exists\,u_{1},\dots,u_{k}\in V_{\text{I}},d(u_{i},u_{j})\geq 2r,\,d_{B}(u_{i})\geq r\>\forall\,i,j=1,\dots,k\right\}\!.$
Theorem 3.11 trivially implies the following corollary, which we state without proof.
Corollary 3.14.
Let $(f,\Lambda)$ be an eigenpair satisfying eq. (18) and assume that $f$ induces $k$ nodal domains. Then:
$\Lambda\geq\frac{1}{R_{k}}\,.$
From this corollary it follows immediately that $\Lambda=1/r_{k}$ with $r_{k}\leq R_{k}$. We proceed now by looking at upper bounds for $\Lambda_{k}$. Recalling that the symbol $\Lambda_{k}$ is used to denote any accumulation point of the sequence $\lambda_{k}^{\frac{1}{p}}(\Delta_{p})$, we have the following Proposition relating $p$-Laplace eigenvalues and packing radii.
Proposition 3.15.
Let $\lambda_{k}(\Delta_{p})$ be the $p$-Laplacian $k$-th variational eigenvalue on a graph $\mathcal{G}$ and let $R_{k}$ be the $k$-th packing radius of $\mathcal{G}$. Then,
$\limsup_{p\rightarrow\infty}\lambda_{k}(\Delta_{p})^{\frac{1}{p}}\leq\frac{1}{R_{k}}\,.$
Proof 3.16.
Let $u_{1},\dots,u_{k}$ be $k$ internal nodes of $\mathcal{G}$ as in Definition 3.13, i.e.:
$d(u_{i},u_{j})\geq 2R_{k}\,,\qquad d_{B}(u_{i})\geq R_{k}\quad\forall i,j=1,\dots,k\,.$
We define $k$ linearly independent functions as:
$f_{i}(u)=\max\{R_{k}-d(u,u_{i}),0\}\,.$
Obviously, the set $A_{k}$ spanned by these functions, $A_{k}=\operatorname{span}\{f_{i}\}_{i=1}^{k}$, has dimension $k$ and also its Krasnoselskii genus $\gamma(A_{k})$ is equal to $k$. Then, by definition, we have:
$\lambda_{k}(\Delta_{p})\leq\max_{f\in A_{k}}\mathcal{R}_{\Delta_{p}}(f)=\max_{f\in A_{k}}\frac{\sum_{(u,v)\in E}\omega_{uv}^{p}|f(u)-f(v)|^{p}}{2\sum_{u\in V}|f(u)|^{p}}\\ =\frac{\sum_{(u,v)\in E}\omega_{uv}^{p}\big{|}\sum_{i=1}^{k}\alpha_{i}f_{i}(u)-\sum_{j=1}^{k}\alpha_{j}f_{j}(v)\big{|}^{p}}{2\sum_{m=1}^{k}\left(\sum_{d(u,u_{m})<R_{k}}|\alpha_{m}f_{m}(u)|^{p}\right)}\,,$
where we recall the factor $1/2$ is in the definition of the $p$-norm of $\nabla_{\omega}f$. If $u$ and $v$ are such that both $d(u,u_{i})$ and $d(v,u_{j})$ are smaller than $R_{k}$, then:
$\omega_{uv}|f(u)-f(v)|$ $=\omega_{uv}|\alpha_{i}||f_{i}(u)-f_{i}(v)\big{|}=\omega_{uv}|\alpha_{i}||d(u,u_{i})-d(v,u_{i})\big{|}$
$\leq\omega_{uv}|\alpha_{i}|d(u,v)\leq\omega_{uv}|\alpha_{i}|\frac{1}{\omega_{uv}}\leq|\alpha_{i}|\,.$
Instead, if $d(u,u_{i})<R_{k}$ and $d(v,u_{j})<R_{k}$ with $i\neq j$, then:
$\omega_{uv}|f(u)-f(v)|$ $=\omega_{uv}|\alpha_{i}\,f_{i}(u)-\alpha_{j}\,f_{j}(v)|$
$\leq\omega_{uv}\,\max\{|\alpha_{i}|,\,|\alpha_{j}|\}\,\big{(}R_{k}-d(u,u_{i})+R_{k}-d(v,u_{j})\big{)}$
$\leq\omega_{uv}\,\max_{l}\{|\alpha_{l}|\}\big{(}2R_{k}-d(u_{i},u_{j})+d(u,v)\big{)}$
$\leq\omega_{uv}\,\max_{l}\{|\alpha_{l}|\}\big{(}2R_{k}-2R_{k}+d(u,v)\big{)}$
$\leq\omega_{uv}\,\max_{l}\{|\alpha_{l}|\}d(u,v)$
$\leq\max_{l}\{|\alpha_{l}|\}\,.$
Lastly, if $d(u,u_{i})<R_{k}$ and $d(v,u_{j})\geq R_{k}$ for all $j$, we have:
$\omega_{uv}|f(u)-f(v)|=$ $\omega_{uv}|\alpha_{i}|\big{(}R_{k}-d(u,u_{i})\big{)}$
$=\omega_{uv}|\alpha_{i}|\big{(}R_{k}-d(u,u_{i})+d(u,v)-d(u,v)\big{)}$
$\leq\omega_{uv}|\alpha_{i}|\big{(}d(u,v)+R_{k}-d(v,u_{i})\big{)}$
$\leq\omega_{uv}|\alpha_{i}|d(u,v)$
$\leq|\alpha_{i}|\,.$
Using the above inequalities in the expression for $\lambda_{k}$, we can write:
$\frac{\sum_{(u,v)\in E}\omega_{uv}^{p}\big{|}\sum\alpha_{i}f_{i}(u)-\sum\alpha_{i}f_{i}(v)\big{|}^{p}}{2\sum_{m=1}^{k}\big{(}\sum_{d(u,u_{m})<R_{k}}|\alpha_{m}f_{m}(u)|^{p}\big{)}}\leq\frac{\sum_{(u,v)\in E}\max_{i}|\alpha_{i}|^{p}}{2\sum_{m=1}^{k}\big{(}\sum_{d(u,u_{m})<R_{k}}|\alpha_{m}f_{m}(u)|^{p}\big{)}}\,.$
Passing to the limit we arrive at:
$\limsup_{p\rightarrow\infty}\lambda_{k}^{\frac{1}{p}}(\Delta_{p})$ $\leq\limsup_{p\rightarrow\infty}\bigg{(}\frac{\sum_{(u,v)\in E}\max_{i}|\alpha_{i}|^{p}}{2\sum_{m=1}^{k}\big{(}\sum_{d(u,u_{m})<R_{k}}|\alpha_{m}f_{m}(u)|^{p}\big{)}}\bigg{)}^{\frac{1}{p}}$
$=\frac{\max_{i}|\alpha_{i}|}{\max_{m,u}|\alpha_{m}f_{m}(u)|}\,.$
The thesis is obtained by observing that $f_{i}(u_{i})=R_{k}$ and thus $\max_{m,u}|\alpha_{m}f(u)|=R_{k}\max_{m}|\alpha_{m}|$.
Next, we show the main theorem of this section, which provides a geometrical characterization of the first two variational eigenvalues of the $p$-Laplacian operator showing that they are exactly equal to the reciprocal of the first two packing radii.
Theorem 3.17.
Assume $\mathcal{G}$ to be a connected graph and $\Lambda_{i}:=\lim_{p_{j}\rightarrow\infty}\lambda_{i}(\Delta_{p_{j}})^{\frac{1}{p_{j}}}$, $i=1,2$. Then
$\Lambda_{i}=\frac{1}{R_{i}}\quad i=1,2\,.$
Proof 3.18.
The proof for the first eigenvalue ($i=1$) follows easily from Proposition 3.15, Corollary 3.9, and the observation that $\lambda_{1}(\Delta_{p})=\min_{f\in\mathcal{H}_{0}(V)}\mathcal{R}_{p}(f)$.
For the second eigenvalue ($i=2$) we proceed as follows. Proposition 3.15 ensures that $\Lambda_{2}\leq\frac{1}{R_{2}}$. Then we can consider a subsequence indexed by $p_{h}$ of convergent eigenpairs, i.e.:
$\big{(}f_{2}(\Delta_{p_{h}}),\lambda_{2}(\Delta_{p_{h}})\big{)}\rightarrow(f,\Lambda_{2})\,.$
From [43, 16], we know that, for any $p_{h}$, $f_{2}(\Delta_{p_{h}})$ has at least two nodal domains, which we denote by $A_{p_{h}}$ for $u\in V_{\text{I}}$ such that $f_{2}(\Delta_{p_{h}})(u)>0$, and $B_{p_{h}}$ for $u\in V_{\text{I}}$ such that $f_{2}(\Delta_{p_{h}})(u)<0$. Then the sets $A=\cap_{n}\cup_{p_{h}>n}A_{p_{h}}$ and $B=\cap_{n}\cup_{p_{h}>n}B_{p_{h}}$ are both non empty and such that $f(u)\geq 0$ for any $u\in A$ and $f(u)\leq 0$ for any $u\in B$. If by contradiction $\{u\>|\;f(u)>0\}=\emptyset$, since $f\neq 0$, there has to exist a node $u$ with $f(u)=0$ that is connected to a node $v\sim u$ such that $f(v)<0$ that means:
$\Delta_{\infty}f(u)=\|\left\{\nabla_{\omega}f\right\}^{-}(u)\|_{\infty}-\|\left\{\nabla_{\omega}f\right\}^{+}(u)\|_{\infty}=\|\left\{\nabla_{\omega}f\right\}^{-}(u)\|_{\infty}>0\,.$
But this is an absurd because $f$ has to satisfy eq. (18), implying that $f$ must induce at least two nodal domains. Then, thanks to Corollary 3.14, we get:
$\Lambda_{2}=\frac{1}{r_{2}}\geq\frac{1}{R_{2}}\,,$
which concludes the proof.
Finally, the following noteworthy Corollary provides lower and upper bounds to the $k$-th variational $\infty$-eigenvalues in terms of packing radii and nodal domains of the corresponding $k$-th eigenfunctions. The proof is a trivial consequence of Proposition 3.15, Corollary 3.14 and Theorem 3.1.
Corollary 3.19.
Let $(f_{k,p_{j}},\lambda_{k}(\delta_{p_{j}}))$ be a sequence of the $k$-th variational $p$-Laplacian eigenpairs. If
$\lim_{j\rightarrow\infty}(f_{k,p_{j}},\lambda_{k}^{\frac{1}{p_{j}}}(\Delta_{p_{j}}))=(f_{k},\Lambda_{k})\,,$
then
$\frac{1}{R_{\mathcal{N}(f_{k})}}\leq\Lambda_{k}\leq\frac{1}{R_{K}}\,,$
where $\mathcal{N}(f_{k})$ is the number or nodal domains induced by $f_{k}$.
Remark 3.20.
It is worthwhile to spend a short remark about the non-boundary case, which is analogous to the homogeneous Neumann case in the continuum setting [22]. In this case, we can think of the Neumann case as a Dirichlet problem defined on a graph $\mathcal{G}$ having a boundary set $B$ formed by nodes $v$ at an infinite distance from any internal node, so that the edge weights $\omega_{uv}$ are zero for all edges such that $u\in V_{\text{I}}$ and $v\in B$. Then it is clear that in this case $\Lambda_{1}=0$ and
$\Lambda_{2}=\frac{1}{R_{2}}:=1/\sup\{r\mathrm{\;\;s.t.\;\;}\exists v_{1},v_{2}\;\>\mathrm{s.t.}\;\>d(v_{1},v_{2})\geq 2r\}\,,$
i.e. half the diameter of the graph. Moreover, for the higher eigenvalues, similarly to Proposition 3.15, we can state that:
$\Lambda_{k}^{-1}\geq\sup\{r>0:\;\exists\;v_{1},\ldots,v_{k}\in V_{\text{I}}:\;d(v_{i},v_{j})\geq 2r\;\>\forall i,j=1,\dots,k\}\,.$
We conclude this section by producing some examples that show that the solution of (18) is not always well defined and that not any solution of (18) is also the limit of $p$-Laplacian eigenfunctions. Moreover we show that, even if $d_{B}$ always realizes the minimum of $\mathcal{R}_{\infty}$, it may not satisfy the limit eigenvalue equation (18).
[…]
Figure 1: Graph of Example 3.21.
Example 3.21.
Consider the graph of figure 1. Easy computations show that both functions $f$ and $g$ given by:
$f=\left(1,\frac{2}{3},\frac{1}{3},\frac{2}{3},1\right)\qquad\mbox{and}\qquad g=\left(1,\frac{2}{3},\frac{1}{3},\frac{2}{3},\frac{4}{9}\right)$
satisfy eq. (18) with $\Lambda=1$. Nevertheless, $g$ cannot be the limit of first eigenfunctions of the $p$-Laplacian, which, because of their uniqueness and the symmetry of the above graph, must be symmetric for any $p$.
[…]
Figure 2: Graph of Example 3.22.
Example 3.22.
Here we consider the distance function on the graph of Figure 2:
$d_{B}(u_{1},u_{2})=\left(\frac{1}{3},\frac{1}{2}\right)\,.$
Then
$\Delta_{\infty}d_{B}(u_{1})=1-2\Big{(}\frac{1}{2}-\frac{1}{3}\Big{)}=1-\frac{1}{3}\gneq 0$
and
$\|\left\{\nabla_{\omega}d_{B}\right\}(u_{1})\|_{\infty}-\Lambda_{1}d_{B}(u_{1})=1-\frac{2}{3}\gneq 0\,.$
Thus $d_{B}$ does not satisfy the $\infty$-limit eigenvalue equation (18), meaning that it can not be the limit of the first eigenfunctions of the $p$-Laplacian.
4 The generalized $\infty$-eigenvalue problem
In this section we discuss a different approach to the $\infty$-Laplacian eigenvalue problem. It is based on the definition of the generalized critical points of a non differentiable functional. The same approach has been deeply investiagted in the study of the graph $1$-Laplacian eigenpairs [11, 27]. More in detail, in the first paragraph, we first recall the definition of the generalized eigenpairs and the corresponding $\infty$-variational eigenvalues. Then, by means of the subgradients of the convex functions $f\mapsto\|f\|_{\infty},\|\nabla_{\omega}f\|_{\infty}$, we establish an explicit algebraic equation satisfied by a generalized eigenpair. In the subsequent paragraphs we prove a topological characterization of the geometrical $\infty$-eigenpairs similar to the one proved in Proposition 3.5. This characterization allows us to compare the two different formulations of the $\infty$-eignepairs. In particular, we prove that any solution of the equation in Theorem 3.1 is a generalized eigenpair, while the generalized eigenpairs solve the limit equation in Theorem 3.1 if we consider a proper subgraph of $\mathcal{G}$.
To generalized $\infty$-eigenpairs are formulated via the generalized “critical points” of the $\infty$-Rayleigh quotient given by:
$\mathcal{R}_{\infty}(f)=\frac{\|\nabla_{\omega}f\|_{\infty}}{\|f\|_{\infty}}\,.$
Indeeed, the $p$-Laplacian eigenpairs can be seen as critical points and values of the functional $\|\nabla_{\omega}f\|_{p}$ on the manifold $\|f\|_{p}=1$, i.e. the points in which the differential of $\|\nabla_{\omega}f\|_{p}$ is normal to $S_{p}=\{f:V_{\text{I}}\rightarrow\mathbb{R}|\;\|f\|_{p}=1\}$. Letting $p\rightarrow\infty$, we observe immediately that $\|\nabla_{\omega}f\|_{\infty}$ is not a differentiable operator and $S_{\infty}$ is not a smooth manifold. Nevertheless, following [13], since $\|\nabla_{\omega}f\|_{\infty}$ and $\|f\|_{\infty}$ are convex functions, we can generalize the notion of critical points as follows:
Definition 4.1.
We say that $(f,\Lambda)$ is a generalized $\infty$-eigenpair if and only if:
$0\in\partial\|\nabla_{\omega}f\|_{\infty}\cap\Lambda\partial\|f\|_{\infty}\,.$
where $\partial\|\nabla_{\omega}f\|_{\infty}$ and $\partial\|f\|_{\infty}$ are the subgradients of the functions $(f\mapsto\|\nabla_{\omega}f\|_{\infty})$ and $(f\mapsto\|f\|_{\infty})$, respectively.
From [13], we observe that Definition 4.1 can be considered as the generalized critical point equation for the functional $(f\mapsto\|\nabla_{\omega}f\|_{\infty})$ on $S_{\infty}$ since $\partial\|\nabla_{\omega}f\|_{\infty}$ is a generalization of the differential of $\|\nabla_{\omega}f\|_{p}$ when $p=\infty$, while, from Lemma 4.2 and 4.3 of [13], (see also [40]) the external cone to $S_{\infty}$ in a point $f$, i.e.,
$C_{Ext}(f)=\{\xi:V\rightarrow\mathbb{R}|\>\left\langle\xi,g-f\right\rangle\leq 0\;\forall g\in S_{\infty}\}$
can be characterized as
$C_{Ext}(f)=\mathop{\cup}_{\lambda\geq 0}\lambda\partial\|f\|_{\infty}\,.$
Moreover we point out that, from Theorem 5.8 of [13], it is possible to introduce the family of Krasnoselskii $\infty$-variational eigenvalues. Similar approaches have been used to study the $1$-Laplacian eigenpairs [11, 27], and more recently to study minimizers of $\|\nabla_{\omega}f\|_{\infty}$ in $L^{2}$ and $L^{\infty}$ spaces, [9, 8].
Definition 4.2.
The Krasnoselskii $\infty$-variational eigenvalues $\Lambda_{k}^{\mathcal{K}}$ is given by:
$\Lambda_{k}^{\mathcal{K}}=\min_{A\in\mathcal{F}_{k}(S_{\infty})}\max_{f\in A}\mathcal{R}_{\infty}(f)\,,$
where, recall $\mathcal{F}_{k}(S_{\infty})$ is the family of the closed symmetric subsets of $S_{\infty}$ with Krasnoselskii genus greater or equal than $k$ (see (12)).
In the next theorem we show that, as shown for the limit variational eigenvalues $\{\Lambda_{k}\}$ in Theorem 3.17 and Corollary 3.19, also the Krasnoselskii variational eigenvalues $\Lambda_{k}^{\mathcal{K}}$ can be related to the packing radii, $R_{k}$, (see Definition 3.13).
Theorem 4.3.
Let $\Lambda_{k}^{\mathcal{K}}$ be the Krasnoselskii variational eigenvalues defined in Definition 4.2 and $R_{k}$ be the $k$-th packing radius defined in Definition 3.13. Then,
$\Lambda_{k}^{\mathcal{K}}\leq\frac{1}{R_{k}}\qquad\forall k=1,\dots,|V|$
and equality holds for $k=1,2$.
Proof 4.4.
For the first part we can follow the proof of Prop 3.15. Consider $u_{1},\dots,u_{k}$ such that $d(u_{i},u_{j})\geq 2r,\;d(u_{i},B)\geq r\;\>\forall i,j=1,\dots,k$ and define the $k$ linearly independent functions
$f_{j}(u)=\max\{R_{k}-d(u,u_{j}),0\}\,,\quad j=1,\dots,k\,.$
The set $A_{k}=span\{f_{j}\}_{j=1}^{k}$ has Krasnoselskii genus equal to $k$, $\gamma(A_{k})=k$, and thus:
$\Lambda_{k}^{\mathcal{K}}\leq\max_{f\in A_{k}}\mathcal{R}_{\infty}(f)\,,$
from which the subsequent inequality readily follows:
$\max_{f\in A_{k}}\mathcal{R}_{\infty}(f)\leq\frac{1}{R_{k}}\,.$
Thus, because of Corollary 3.9, we are left to prove that $\Lambda_{2}^{\mathcal{K}}\geq R_{2}^{-1}$. To do this, we first recall that $\mathcal{F}_{2}(S_{\infty})$ is formed by the closed and symmetric subsets of $\mathcal{H}_{0}(V)\cap S_{\infty}$ As a consequence, we observe that any $A\in\mathcal{F}_{2}(S_{\infty})$ necessarily contains a symmetric, closed, and connected subset of $S_{\infty}$. Then, given the function
$\psi:$ $S_{\infty}$$\rightarrow \displaystyle\mathbb{R}$
$f$ $\mapsto$ $\|f^{+}\|_{\infty}-\|f^{-}\|_{\infty}\,,$
for any $A\in\mathcal{F}_{2}(S_{\infty})$ there esists $f_{A}\in A$ such that $\psi(f_{A})=0$. In other words, there exist $u^{+},u^{-}\in V$ such that
$\|f\|_{\infty}=f_{A}(u^{+})=-f_{A}(u^{-})\,.$
To conclude, from (23), we observe that
$d(u^{\pm},B)\geq\frac{\|f_{A}\|_{\infty}}{\|\nabla_{\omega}f_{A}\|_{\infty}}\,,\qquad d(u^{+},u^{-})\geq 2\frac{\|f_{A}\|_{\infty}}{\|\nabla_{\omega}f_{A}\|_{\infty}}\,,$
i.e. :
$\frac{\|f_{A}\|_{\infty}}{\|\nabla_{\omega}f_{A}\|_{\infty}}\leq R_{2}\qquad\Rightarrow\qquad\Lambda_{2}^{\mathcal{K}}\geq\min_{A\in\mathcal{F}_{2}}\frac{\|\nabla_{\omega}f_{A}\|_{\infty}}{\|f_{A}\|_{\infty}}\geq\frac{1}{R_{2}}\,.$
Next, we study the structure of the two sets $\partial\|\nabla_{\omega}f\|_{\infty}$ and $\partial\|f\|_{\infty}$. As in remark 2.6, we think about both $f\mapsto\|f\|_{\infty}$ and $f\mapsto\|\nabla_{\omega}f\|_{\infty}$ as functions from $\mathcal{H}(V_{\text{I}})$ to $\mathbb{R}$. Then, as done in [9], because of homogeneity, it is easy to derive the following characterization, (see also [10]):
$\partial\big{(}f\mapsto\|f\|_{\infty}\big{)}=\{\xi:V_{\text{I}}\to\mathbb{R}\,|\,\|g\|_{\infty}\geq\left\langle\xi,g\right\rangle\,\forall g:V_{\text{I}}\to\mathbb{R},\;\>\|f\|_{\infty}=\left\langle\xi,f\right\rangle\}\,,$
i.e.
(26) $\partial\|f\|_{\infty}:=\left\{\xi\;\bigg{|}\begin{array}[]{lr}\|\xi\|_{1,V}=1,\;\xi(u)=0\;\;\forall u\in V_{\text{I}}\setminus V_{\text{max}}(f)\,,\\[3.99994pt] |\,\xi(u)|\,|f(u)|=\xi(u)f(u)\;\;\forall u\in V_{\text{max}}(f)\end{array}\right\}\,.$
Moreover, we can use the subdifferential chain rule for linear transformations $\Big{(}\partial\big{(}x\mapsto\phi(Ax)\big{)}=A^{T}\partial\big{(}y\mapsto\phi(y)\big{)}|_{y=Ax}\Big{)}$ to characterize $\partial(f\mapsto\|\nabla_{\omega}f\|_{\infty})$ [40]:
(27) $\partial\|\nabla_{\omega}f\|_{\infty}:=\left\{-\operatorname{div}\Xi\,\bigg{|}\!\!\begin{array}[]{lr}\|\Xi\|_{1,E}=1,\;\Xi(u,v)=0\;\;\forall(u,v)\in E\setminus E_{\text{max}}(f)\,,\\[3.99994pt] |\,\Xi(u,v)\,||\nabla_{\omega}f(u,v)|=\Xi(u,v)\nabla_{\omega}f(u,v)\;\forall(u,v)\in E_{\text{max}}(f)\end{array}\!\!\!\right\}$
where the divergence operator in defined as in Remark 2.6, and we used the edge and node norms defined in (8). Then we can give the following definition of generalized $\infty$-eigenpair.
Definition 4.5.
$(f,\Lambda)$ is a generalized $\infty$-eigenpair if and only if there exist $\xi\in\partial\|f\|_{\infty}$ and $\Xi$ with $-\operatorname{div}(\Xi)\in\partial\|\nabla_{\omega}f\|_{\infty}$ such that
$-\operatorname{div}\,\Xi(u)=\Lambda\xi(u)\quad\forall u\in V_{\text{I}}\,.$
Note that, putting together the above definition of generalized eigenpair and the characterization of the subgradients $\partial\|f\|_{\infty}$ and $\partial\|\nabla_{\omega}f\|_{\infty}$ (eqs. (26) and (27)), we have the following Proposition, whose proof is immediate.
Proposition 4.6.
$(f,\Lambda)$ with $f\in\mathcal{H}_{0}(V)$, is a generalized $\infty$-eigenpair if and only if there exist $\xi:V_{\text{I}}\rightarrow\mathbb{R}$ and $\Xi:E\rightarrow\mathbb{R}$ such that
(28) $\begin{cases}-\operatorname{div}\Xi(u)=\Lambda\xi(u)\quad&\forall u\in V_{\text{I}}\\ \|\Xi\|_{1,E}=1\\ \|\xi\|_{1,V}=1\\ |f(u)|=\|f\|_{\infty}\quad&if\;\xi(u)\neq 0\\ |\nabla_{\omega}f(u,v)|=\|\nabla_{\omega}f\|_{\infty}\quad&if\;\Xi(u,v)\neq 0\\ \mathrm{sign}\big{(}\Xi(u,v)\big{)}=\mathrm{sign}\big{(}\nabla_{\omega}f(u,v)\big{)}\qquad&if\;\Xi(u,v)\neq 0\\ \mathrm{sign\big{(}\xi(u)\big{)}}=\mathrm{sign\big{(}f(u)\big{)}}\quad&if\;\xi(u)\neq 0\end{cases}\,.$
Moreover, up to redefining $\Xi(u,v)=\Big{(}\Xi(u,v)-\Xi(v,u)\Big{)}/2$, we can assume $\Xi(u,v)=-\Xi(v,u)$.
Remark 4.7.
We would like to observe that since $\mathcal{R}_{\infty}$ is a locally Lipschitz function of $\mathbb{R}^{n}\setminus\{0\}$, the notion of critical point can also be generalized considering the Clarke subderivative $\partial^{Cl}\mathcal{R}_{\infty}$, see [14], i.e. $f$ is a Clarke $\infty$-eigenpairs iff
(29) $0\in\partial^{Cl}\mathcal{R}_{\infty}(f)\,.$
From [14, 27] we know that:
(30) $\partial^{Cl}\mathcal{R}_{\infty}(f)\subseteq\frac{\partial\|\nabla_{\omega}f\|_{\infty}\|f\|_{\infty}-\partial\|f\|_{\infty}\|\nabla_{\omega}f\|_{\infty}}{\|f\|^{2}_{\infty}}\,,$
which shows that the notion of generalized $\infty$-eigenpair (Definition (4.5)) extends the notion of Clarke-eigenpair. We refer to the work in [45] for an example where the set of the Clarke eigenvalues is strictly included in the set of the generalized eigenvalues. Nevertheless, Definition (4.5) provides some practical advantages with respect to the one (29) since, differently from $\partial^{Cl}\mathcal{R}_{\infty}(f)$, both $\partial\|f\|_{\infty}$ and $\partial\|\nabla_{\omega}f\|_{\infty}$ can be explicitly identified.
4.1 Topological characterization
In this section we discuss a geometrical characterization of the generalized $\infty$-eigenfunctions similar to the one proved in Proposition 3.5. In both these characterizations, given an eigenpair, $(f,\lambda)$, we prove the existence of “good” paths that connect points in $V_{\text{max}}(f)\cup B$ and whose length matches the value of the eigenvalue. However, differently from Proposition 3.5, where we proved the existence of a “good” path $\Gamma$ for any extremal point $v\in V_{\text{max}}(f)$, in the case of generalized $\infty$-critical points, extremal points that do not correspond to any “good” path may indeed exist. Moreover, contrary to the case of limit $\infty$-eigenpairs where the existence of such “good” paths was only a necessary condition for the existence of an eigenpair, in the case of generalized eigenpairs, the existence of “good” paths is also a sufficient condition.
Proposition 4.8.
The pair $(f,\Lambda)$ is a generalized $\infty$-eigenpair with $f$ not a constant function if and only if there exists a path $\Gamma=\{(u_{i},u_{i+1})\}_{i=1}^{n-1}$ such that:
  1. 1.
    $u_{i}\in V_{\text{max}}(f)\cup B\quad i=1,n$;
  2. 2.
    $(u_{i},u_{i+1})\in E_{\text{max}}(f)$ and $f(u_{i})>f(u_{i+1})$ for all $i=1,\dots,n$;
  3. 3.
    assuming without loss of generality that $f(u_{1})>0$ , if $u_{n}\in B$ then $\Lambda=\frac{1}{\mathrm{length}(\Gamma)}$, if $u_{n}\notin B$ then $f(u_{n})=-f(u_{1})$ and $\Lambda=\frac{2}{\mathrm{length}(\Gamma)}$. Moreover
    $\frac{1}{\Lambda}=\min\bigg{\{}\min_{\{v|f(v)=-\|f\|_{\infty}\}}\frac{d(u_{1},v)}{2},d_{B}(u_{1})\bigg{\}}\,.$
Proof 4.9.
Assume that $(f,\Lambda)$ is a generalized $\infty$-eigenpair and let $(\xi,\Xi)$ be the two subgradients as in Definition (28), i.e., such that
(31) $-\operatorname{div}(\Xi)=\Lambda\xi\,,$
and $\Xi(u,v)=-\Xi(v,u)$. Let $\xi(u_{1})\neq 0$ and without loss of generality assume $f(u_{1})>0$. From eq. (28) there has to exist an edge $(u_{1},u_{2})\in E_{\text{max}}(f)$ such that $\Xi(u_{1},u_{2})\neq 0$, i.e. $f(u_{2})=f(u_{1})-\|\nabla_{\omega}f\|_{\infty}/\omega_{u_{1}u_{2}}<f(u_{1})$. Let us focus now on the second node $u_{2}$. Since $f(u_{2})<f(u_{1})$, if $u_{2}\notin B$ and $f(u_{2})\neq-f(u_{1})$, we have that $\xi(u_{2})=0$. Moreover $\Xi(u_{1},u_{2})<0$ and $\Xi$ has to satisfy (31), i.e.:
$-\operatorname{div}(\Xi)(u_{2})=\sum_{v\sim u_{2}}\omega_{vu_{2}}\Xi(v,u_{2})=\Lambda\xi(u_{2})=0\,,$
hence, there must exist an edge $(u_{3},u_{2})\in E_{\text{max}}(f)$ such that $\Xi(u_{2},u_{3})>0\,,$ i.e., $f(u_{3})=f(u_{2})-\|\nabla_{\omega}f\|_{\infty}/\omega_{u_{2}u_{3}}<f(u_{2})$ Iterating the above argument, we can define a path $\Gamma=\{(u_{i},u_{i+1})\}_{i=1}^{n-1}$ such that $u_{n}\in B\cup V_{\text{max}}(f)$, $(u_{i},u_{i+1})\in E_{\text{max}}(f)$ and $f(u_{i})>f(u_{i+1})$. Furthermore, as in the proof of Proposition 3.5, it is easy to see that, given a function $f$ and a path $\Gamma$ that satisfies the first two items of the thesis, necessarily we have that, if $u_{n}\in B$, $\Lambda=1/\mathrm{length}(\Gamma)$. On the other hand, if $f(u_{n})=-f(u_{1})$, $\Lambda=2/\mathrm{length}(\Gamma)$ and
$\frac{1}{\Lambda}=\min\bigg{\{}\min_{f(v)=-\|f\|_{\infty}}\frac{d(u_{0},v)}{2},d_{B}(u_{0})\bigg{\}}\,.$
Indeeed, proceeding by contraddiction, if one of the last equalities was not true then an edge $(v_{1},v_{2})$ with $|\nabla_{\omega}f(v_{1},v_{2})|>\|\nabla_{\omega}f\|_{\infty}$ would exist.
To prove the opposite inclusion, we assume that, given $(f,\Lambda)$, there exists a path $\Gamma=\{(u_{i},u_{i+1})\}_{i=1}^{n-1}$ with $f(u_{1})>0$ that satisfies the thesis and look at the case $u_{n}\in\{v\,|\,f(v)=-f(u_{1})\}$ with $\Lambda=2/d(u_{1},u_{n})=2/\mathrm{length}(\Gamma)<d_{B}(u_{1})$ (the other case can be proved analogously). We aim at defining two functions $\Xi$ and $\xi$ that satisfy the sufficient conditions in Proposition 4.6. Define the two functions as follows:
$\Xi(u,v):=\frac{\delta_{\Gamma}(u,v)}{\omega_{uv}}\>\frac{\mathrm{sign}\big{(}\nabla_{\omega}f(u,v)\big{)}}{\mathrm{length}(\Gamma)}$
$\xi(v):=\frac{\delta_{u_{1}}(v)\mathrm{sign}\big{(}f(u_{1})\big{)}+\delta_{u_{n}}(v)\mathrm{sign}\big{(}f(u_{n})\big{)}}{2}\,,$
where $\delta_{\Gamma}$ $\delta_{u_{i}}$ are delta functions, i.e., $\delta_{\Gamma}(u,v)=1$ if $(u,v)\in\Gamma$ and $\delta_{\Gamma}(u,v)=0$ otherwise, with analogous definition for $\delta_{u_{i}}$. Then, since $\xi$ and $-\operatorname{div}\Xi$ belong respectively to the two subgradients $\partial\|f\|_{\infty}$ and $\partial\|\nabla_{\omega}f\|_{\infty}$, an easy calculations shows that for any node $v\not\in\Gamma$ as well as for any node $v\in\Gamma$, with $v\neq u_{1}$ and $v\neq u_{n}$, we have:
$-\operatorname{div}\Xi(v)=0=\Lambda\,\xi(v)\,.$
Finally, in the case $v=u_{1}$ (or analogously $v=u_{n}$), we have
$-\operatorname{div}\Xi(u_{1})=\frac{\omega_{u_{1}u_{2}}}{2}\Big{(}\Xi(u_{2},u_{1})-\Xi(u_{1},u_{2})\Big{)}=\frac{\mathrm{sign}\big{(}(\nabla_{\omega}f(u_{2},u_{1})\big{)}}{\mathrm{length}(\Gamma)}=\frac{\Lambda}{2}=\Lambda\xi(u_{1})\,,$
where we have used the fact that $\mathrm{sign}\big{(}\nabla_{\omega}f(u_{2},u_{1})\big{)}=\mathrm{sign}\big{(}f(u_{1})-f(u_{2})\big{)}=1$, which follows by the hypothesis $f(u_{1})>f(u_{2})$. This concludes the proof.
As a corollary of Proposition 4.8 and Proposition 3.5 it is easy to prove that any eigenpair that satisfies the limit eigenvalue equation (18) is also a generalized $\infty$-eigenpair.
Corollary 4.10.
Let $(f,\Lambda)$ satisfy the limit eigenvalue equation (18) with $f$ not constant, then $(f,\Lambda)$ is also a generalized $\infty$-eigenpair according to Definition 4.5.
Proof 4.11.
The proof is a consequence of the fact that, from Propositon 3.5, for any eigenpair that satisfies the limit eigenvalue equation (18), there exists a path $\Gamma$ that satisfies the hypotheses of Proposition 4.8.
Next we deal with the opposite problem and ask ourselves if any generalized $\infty$-eigenvalue (see Definition 4.1) can be associated to an eigenfunction that solves (18). As we will see in Example 4.17, the answer is negative in general. Nevertheless, in Theorem 4.15 we will see that the statement can always be proved to hold up to a subgraph of $\mathcal{G}$. Before we prove this statement and discuss the properties subgraph, we observe that the $\infty$-eigenvalue problem can be reformualted in terms of a constrained weighted linear Laplacian eigenvalue problem. Indeed, from the characterizations of the subgradient in (27) and (26), it is possible to reformulate system (28) as in the following proposition.
Proposition 4.12.
The pair $(f,\Lambda)$ is a generalized $\infty$-eigenpair if and only if there exist two admissible densities $\nu:V_{\text{I}}\rightarrow\mathbb{R}_{+}$, and $\mu:E\rightarrow\mathbb{R}_{+}$, with $\mu_{uv}=\mu_{vu}$ such that:
(32) $\begin{cases}-\operatorname{div}(\mu\odot\nabla_{\omega}f)(u)=\Lambda\nu_{u}f(u)\quad&\forall u\in V_{\text{I}}\\ |\nabla_{\omega}f(u,v)|=\|\nabla_{\omega}f(u,v)\|_{\infty}\quad&\text{if}\;\mu_{uv}>0\\ |f(u)|=\|f(u)\|_{\infty}\quad&\text{if}\;\nu_{u}>0\\ \|\mu\odot\nabla_{\omega}f\|_{1,E}=1\\ \|\nu\odot f\|_{1,V}=1\\ f(u)=0\quad&if\;u\in B\end{cases}\,.$
Proof 4.13.
Straightforward substitution into equation (28) shows that the quantities
$\nu_{u}:=\frac{|\xi(u)|}{2\|f\|_{\infty}}\qquad\mu:=\frac{|\Xi(u,v)|+|\Xi(v,u)|}{\|\nabla_{\omega}f\|_{\infty}}$
are the desired admissible densities. The inverse implication follows by inverting the above definitions of $\nu$ and $\mu$.
Remark 4.14.
We remark the an analogue linearization of the graph $p$-Laplacian eigenvalue problem, when $p\in(2,\infty)$, has been used in [17] to compute $p$-eigenpairs as limits of sequences of linear eigenpairs. Thus, in a similar way, the linearization of the $\infty$-eigenproblem in Proposition 4.12 could be the starting point to investigate numerical methods aimed at computing generalized $\infty$-eigenpairs, we refer to [15] for a preliminary investigation. On the other hand, we refer to [5] for a numerical investigation of the solutions of the $\infty$-limit eigenvalue equation.
To continue towards the sought result we need some definitions. Let $(f,\Lambda)$ denote an $\infty$-eigenpair and assume $(\mu,\nu)$ to satisfy the condition of Proposition 4.12. We say that a node $u\in V$ is supported by $\mu$, ($u\in V_{\mu}$), if there exists an edge $(u,v)\in E$ such that $\mu_{uv}>0$. Observe that if $u\in V_{\text{I}}$ but $u\not\in V_{\mu}$, then necessarily $\nu_{u}=0$ (recall that if $\nu_{u}\neq 0$, $|f(u)|=\|f\|_{\infty}$). In particular, we write $u\in supp(V)$, if there exist at least one $(\mu,\nu)$ as in Proposition 4.12, such that $u\in V_{\mu}$. Now we can prove that any generalized $\infty$-eigenpair can be regarded as a limit $\infty$-eigenpair up to considering the subraph of $\mathcal{G}$ induced by $\mbox{supp}(V)$.
Theorem 4.15.
Assume $(f,\Lambda)$ to be an $\infty$-eigenpair as in Definition 4.1. If $u\in\mbox{supp}(V)\cap V_{\text{I}}$, then $f$ satisfies the limit eigenvalue equation (18) in $u$.
Proof 4.16.
Let $u\in\mbox{supp}(V)\cap V_{\text{I}}$ and let $(\mu,\nu)$ be a pair of admissible densities, such that $u\in V_{\mu}$. Assuming $f(u)>0$, the weighted eigenvalue equation $-\operatorname{div}(\mu\nabla_{\omega}f)(u)=\Lambda\nu_{u}f(u)$ can be written as:
(33) $-\|\nabla_{\omega}f\|_{\infty}\sum_{v\sim u}\mu_{uv}\,\omega_{uv}\mathrm{sign}\big{(}\nabla_{\omega}f(u,v)\big{)}=\Lambda\nu_{u}f(u)=\frac{\|\nabla_{\omega}f\|_{\infty}}{\|f\|_{\infty}}\nu_{u}f(u)\,.$
We first consider the case $f(u)<\|f\|_{\infty}$. Then $\nu_{u}=0$, which yields $\|\left\{\nabla_{\omega}f\right\}^{-}(u)\|_{\infty}=\|\left\{\nabla_{\omega}f\right\}^{+}(u)\|_{\infty}$ i.e. $\Delta_{\infty}f(u)=0$, and:
$\|\left\{\nabla_{\omega}f\right\}(u)\|_{\infty}-\Lambda f(u)=\|\nabla_{\omega}f\|_{\infty}-\frac{\|\nabla_{\omega}f\|_{\infty}}{\|f\|_{\infty}}f(u)>0\,.$
If instead $f(u)=\|f\|_{\infty}$, for any $v\sim u$ we get $\nabla_{\omega}f(u,v)\leq 0$ and thus $\nu_{u}\neq 0$ (otherwise, by hypotheses, (33) is not satisfied). Then, eq. (33) reads:
$\|\nabla_{\omega}f\|_{\infty}\Big{(}\sum_{v\sim u}\mu_{uv}\omega_{uv}\Big{)}=\frac{\|\nabla_{\omega}f\|_{\infty}}{\|f\|_{\infty}}\nu_{u}\|f\|_{\infty}\,,$
[…]
$\begin{cases}\|\left\{\nabla_{\omega}f\right\}(u)\|_{\infty}=\|\nabla_{\omega}f\|_{\infty}=\Lambda f(u)\,,\\ \Delta_{\infty}f(u)=\|\left\{\nabla_{\omega}f\right\}^{-}(u)\|_{\infty}=\|\nabla_{\omega}f\|_{\infty}>0\,.\end{cases}$
The proof is concluded by considering the cases $f(u)<0$ and $f(u)=0$, which can be proved analogously.
We conclude this section by showing that there exist $\infty$-eigenpairs that are generalized critical points of $\mathcal{R}_{\infty}$ but that do not satisfy the $\infty$-eigenvalue equation (18). By the same example, we also prove that there exist generalized $\infty$-eigenvalues concentrated between the first and second variational eigenvalues, which on the contrary never exist for the $p$-Laplacian when $p\in(1,\infty)$ [16].
[…]
Figure 3: A path graph admitting $\infty$-generalized eigenvalues that are not $\infty$-limit eigenvalues
Example 4.17.
Consider the graph of figure 3: Note that the node farthest from the boundary is $u_{2}$ and $d(u_{2},B)=1/2+1/3=5/6$. Then, the pair:
$f_{1}(u_{2},u_{3})=\left(\frac{5}{6},\,\frac{1}{2}\right)\;,\;\Lambda_{1}=\frac{6}{5}$
is an $\infty$-eigenpair with
[…]
However, it is easy to verify that the following are also eigenpairs:
[…]
where $*\in{\left[\frac{1}{6},\frac{1}{2}\right]}$ means that any value in $[1/6,1/2]$ produces an eigenfunction corresponding to the eigenvalue $\Lambda=2$. Note that $\Lambda_{2}=R_{2}=6$ is the second variational eigenvalue while $\Lambda_{1}<\Lambda=2<\Lambda_{2}$. Moreover, it is worth noting that the pair $(f,\Lambda)$ does not satisfy (18). Indeed, to get a solution of (18) $f(u_{2})$ should be determined in such a way as to get:
$\Delta_{\infty}f(u_{2})=0,\quad\|\left\{\nabla_{\omega}f\right\}(u_{2})\|_{\infty}-2f(u_{2})\geq 0$
or
$\Delta_{\infty}f(u_{2})\geq 0,\quad\|\left\{\nabla_{\omega}f\right\}(u_{2})\|_{\infty}-2f(u_{2})=0\,.$
In the first case, $\Delta_{\infty}f(u_{2})=0$ implies:
$f_{2}(u_{2})=\frac{3}{8}\;\Rightarrow\;\|\left\{\nabla_{\omega}f_{2}\right\}(u_{2})\|_{\infty}-2f(u_{2})=\frac{3}{8}-\frac{6}{8}<0\,.$
In the second case, the equality $\|\left\{\nabla_{\omega}f\right\}(u_{2})\|_{\infty}-2f(u_{2})=0$ implies $f_{2}(u_{2})=\frac{3}{10}$ from which we get:
$\Delta_{\infty}f_{2}(u_{2})=\frac{3}{10}-\frac{3}{5}<0\,.$